### v2

# Modeling Human Behavior in a Strategic Network Game with Complex Group Dynamics

Abstract
Human networks greatly impact important societal outcomes, including wealth and health inequality, poverty, and bullying. As such, understanding human networks is critical to learning how to promote favorable societal outcomes. As a step toward better understanding human networks, we compare and contrast several methods for learning, from a small data set, models of human behavior in a strategic network game called the Junior High Game (JHG) [41]. These modeling methods differ with respect to the assumptions they use to parameterize human behavior (behavior vs. community-aware behavior) and the moments they model (mean vs. distribution). Results show that the highest-performing method, called hCAB, models the distribution of human behavior rather than the mean and assumes humans use community-aware behavior rather than behavior matching. When applied to small societies (6-11 individuals), the hCAB model closely mirrors the population dynamics of human groups (with notable differences). Additionally, in a user study, human participants were unable to distinguish individual hCAB agents from other humans, thus illustrating that the hCAB model also produces plausible (individual) human behavior in this strategic network game.
1 Introduction
Human networks, collections of people linked through social connections, influence how individuals exchange resources [26], form alliances [12, 20], and navigate competitive pressures [8, 46]. These networks, which are studied in many disciplines [43], exhibit emergent behaviors related to inequality [6, 11, 37], information flow [9] and social behaviors [40, 10, 45]. While the study of human networks has yielded deep insights [29, 13, 5, 35], understanding the precise decision-making processes that drive human behavior within these networks remains a major challenge [49]. What determines whether an individual cooperates or competes? How do local interactions scale up to produce global patterns of stability, inequality, or conflict? How might an individual change their behavior to encourage preferred social change? Addressing these questions requires computational models that capture the complexities of human decision-making in networked environments.
In human networks, each individual makes decisions based on personal incentives, past experiences, and the social landscape. Sometimes individual incentives align with the collective good, like the invisible hand [42], but other times they are diametrically opposed, like the free rider problem [34]. These individual actions lead to emergent behaviors that shape societal outcomes. Extensive studies have explored emergent properties of human networks [22, 23, 39], investigating, for example, how randomness affects network structures [14], how power distributes among individuals [6], and how self-organization arises within different social [47] and economic settings [1, 48]. These macroscopic perspectives on how human societies function are valuable, yet they often are insufficient for fully modeling human networks and how they evolve. More detailed models of human decision-making are needed to better understand human networks.
To help address this challenge, we consider modeling human behavior in the Junior High Game (JHG) [41], a controlled game environment designed to capture the strategic trade-offs in human networks. In this game, players engage in repeated economic exchanges, choosing whether to give, keep, or take resources from others. Decisions are shaped by social factors such as dyadic relationships [17, 15] triadic (indirect) interactions [18, 33], and group considerations [31, 7, 21]. JHG scenarios produce networks that are dynamic, directed, weighted, and signed, thus providing a rich testbed for studying human behavior in networked settings.
A fundamental challenge in modeling human decision-making in the JHG lies in determining how to parameterize and represent behavior [36]. In this paper, we consider models that make different assumptions about human decision-making. One approach, behavior matching [4], assumes that agents imitate the observed actions of others, while another approach, community-aware behavior [41], considers broader structural and social cues within the network. We also examine how modeling different statistical moments (the mean verses the distribution of behavior) impact model performance. Together, this produces four different algorithms for modeling human behavior in the JHG, which we compare and contrast in small-scale societies (6-11 players).
We evaluate the ability of the resulting modeling algorithms to produce human behavior in the JHG in two ways. First, we compare and contrast the population dynamics of agent societies with those of human groups. Results show that societies of one of these agents, called hCAB, have population dynamics that closely mirror (with some differences) human populations. Second, we conduct a user study to identify the extent to which individual hCAB agents behave like human players. Results of this study show that participants were unable to distinguish hCAB agents from other human players. These findings not only inform the development of models that can simulate human behavior in the JHG, but also offer perspectives on the mechanisms that drive the dynamics of real-world networks.
2 A strategic network game to model human networks
In this paper, we model human networks using the Junior High Game (JHG) [41]. The JHG models key properties of human networks including conflicting interests, a large and flexible strategy space, high interdependence among agents, power asymmetry, and (often) non-convergence [23, 44]. The JHG induces a dynamic (time varying), directed, weighted, and signed network.
We paraphrase Skaggs et al. [41] to review this game: The JHG is played by a set of players $I$ over a series of rounds in which each player $j\in I$ seeks to become popular. Initially, every player is assigned a popularity, which changes over time based on interactions, represented by token allocations, among the players. In every round, each player $j$ allocates $N$ tokens, each of which it can keep, give to another player, or attack another player. Keeping tokens positively impacts $j$’s subsequent popularity, while tokens $j$ gives to player $i$ positively impact $i$’s subsequent popularity (but do nothing for $j$). Finally, when $j$ attacks $i$, it negatively impacts $i$’s subsequent popularity while positively impacting $j$’s. After all players allocate their tokens in a round, the players’ popularities are updated based on the players’ token allocations. A new round then begins.
The amount that a token transaction impacts subsequent popularity depends on the popularity of the player allocating the token. For example, receiving a token from a more popular individual results in a higher increase in popularity than receiving a token from a less popular individual. Thus, players with higher popularity can have higher impact than those with lower popularity.
(a)                                            (b)                                 (c)
Figure 1: Example scenario illustrating various JHG dynamics. (a) Popularity of the players over time. (b) Token allocations in rounds 1-4. Red dashed arrows indicate tokens used to attack while black solid arrows indicate tokens given. (c) Token allocations in rounds 5-6.
To increase understanding of JHG dynamics, consider the contrived example depicted in Figure 1. First, when players exchange (give) tokens with each other, they rise in popularity (e.g., players p2 and p3). Second, when a player gives tokens without receiving as much in return, they tend to decrease in popularity (e.g., p4). Third, when a player keeps its token, it maintains much of its popularity, but does not increase in popularity in the absence of other interactions (e.g., p1 in rounds 1-4). Fourth, attacking those not keeping tokens results in the attacker becoming more popular and the victim becoming less popularity (e.g., p5 attacking p7 and p8 in rounds 1-4). When other players do not keep or retaliate, attacking is the fastest way to gain popularity. It also increases the attacker’s relative popularity in relation to the victim. Fifth, keeping tokens blocks attacks. For example, p6’s attacks on p1 in rounds 1-4 are blocked because p1 is keeping its tokens–p1 experiences no loss, while p6 quickly loses popularity due to not gaining from its token allocations. Finally, attacks that are stronger than the defense (strength is measured as tokens allocated multiplied by player popularity) result in inflicted damage. For example, p5’s attack, combined with p6’s, on p1 in rounds 5-6 is substantially stronger than p1’s defense. Thus, p1 loses popularity and p5 and p6 profit somewhat.
While the contrived example presented above deals primarily with pairwise interactions for illustrative purposes, these dynamics can give rise to complex group behavior and emergent social norms. Frequently, alliances form and dissolve as players join together to help each other, punish those that violate norms, mitigate powerful enemies, and exploit others.
3 Modeling human behavior in the JHG using human data
In this section, we introduce four algorithms for learning models of human behavior in the JHG from a relatively small data set of games played by experienced human players. These algorithms vary based on assumptions encoded in their parameter-based representations and their methods for learning parameterizations. We discuss these design choices in turn.
3.1 Parameter-based models of behavior
We construct algorithms based on two parameter-based models of behavior for playing the JHG: Behavior Matching (TFT) and community-aware behavior (CAB).
Behavior Matching (TFT): In behavior matching, or tit-for-tat (TFT), a player reciprocates the actions that other players directed towards them in the previous round [4]. We consider this model due to its prevalence in prior work. Tit-for-tat and its variants have repeatedly been shown to be effective and robust in repeated prisoner’s dilemmas (e.g., [3, 32, 19]), a common domain for studying when and how humans and other agents cooperate with each other.
Defining TFT in the JHG is not straightforward, as a player may receive more or less in a round than it has the ability to reciprocate. In such cases, an agent must determine who to exclude (when receiving too much) or what to do with extra tokens (when not receiving enough). We have defined a TFT algorithm for the JHG whose behavior is tuned with seven parameters. These parameters impact how the agent allocates tokens in the first round, what it seeks to match (token allocations or influence), and how its deals with excess and shortfalls. Algorithmic details are given in SM-B.
Community-Aware Behavior (CAB): The second parameter-based model we consider is the CAB (Community Aware Behavior) algorithm [41]. While behavior matching assumes humans primarily reciprocate the actions of others, the CAB agent is based primarily on group dynamics. A CAB agent determines token allocations in a round using a three-step process. First, it assigns players to groups based on past token allocations. Second, it chooses a (potentially new) group it would like to form, support, or join. Finally, it selects token allocations designed to build up its selected group, establish itself in the group, and deal with out-of-group threats. CAB behavior is defined by thirty parameters which determine how it identifies existing groups, chooses a group to try to be part of, and allocates tokens to build its group. We use the CAB implementation provided by Skaggs et al. [41].
Assuming that human behavior in the JHG can be modeled using either of these algorithms is most certainly incorrect at some level. However, identifying the extent to which these models can and cannot approximate human behavior can shed valuable insights into human networks.
3.2 Learning parameterizations
To determine how well either parameter-based model (TFT or CAB) can model human behavior in the JHG, we learn parameterizations of each model from a limited amount of human data using two different methods: Particle Swarm Optimization (PSO) [25] and Evolutionary Population Distribution Modeling (EPDM), a methodology proposed in this paper. Both algorithms take as input a parameter-based model and output one or more parameterizations of the model. Let $m$ be the number of parameters in the desired model. Then, $\Theta\in(\theta_{1},\theta_{2},\cdots,\theta_{m})$ is a parameterization of the model in which $\theta_{i}$ is the value of the $i^{\rm th}$ parameter.
Particle Swarm Optimization (PSO): In PSO, a set of particles (i.e., parameterizations) are first randomly generated. Over time, these parameterizations are adapted to both explore the parameter space and converge toward the best parameterization. In our case, the best parameterization is determined by minimizing the error, over all data samples, between the token allocations made by human players and the token allocations made by the model (given the parameterization specified by the particle) in the same situations. In this way, PSO produces a parameterization that describes mean behavior in the data set.
We experimented with two different error (or distance) functions: (1) mean squared error and (2) a custom built function in which the error is based on how well token allocations match across a variety of different properties (defined in SM-D). Because the second function worked best in all cases, we used that error function in all of our experiments.
Evolutionary Population Distribution Modeling (EPDM): Because mean human behavior, when followed by all individuals, does not necessarily produce human-like population dynamics or realistic individual behavior, we introduce a second method, called EPDM, for modeling human behavior from data. EPDM attempts to model the distribution of human behavior, rather than mean behavior.
The EPDM algorithm, outlined in Algorithm 1, receives as input a parameter-based behavior model $\mathcal{M}$ (TFT or CAB) and a set of player games $\Gamma$ (from the training data). Each player game $\gamma\in\Gamma$ specifies the token allocations made by a human player in each round $t$ of a game, along with the game state available to the player at the beginning of each round (specifically, ${\mathcal{P}}(t)$, ${\mathcal{I}}(t)$, and $X(t)$). The algorithm outputs a set of $N$ (we used 100) parameterizations of $\mathcal{M}$ that best estimate the distribution of player strategies in $\Gamma$.
Algorithm 1 EPDM – Computes $N$ parameterizations to estimate a distribution of strategies
  Input: $\mathcal{M}$ (parameter-based model), $\Gamma$ (player games), and $G$ (# generations)
  Initialize: Set $g=0$ and randomly generate $\Pi(0)$
  while $g<G$ do
     (1) $\forall\pi\in\Pi(g)$, compute $\Gamma^{\pi}(g)$ (player games in which $\pi$ is a top performer)
     (2) $\Phi(g)=$ FindCoreSet$(\Gamma,\Pi(g),\forall{\pi\in\Pi(g)},~{}\Gamma^{\pi}(g))$
     (3) Construct $\Pi(g+1)$:
           - $\Pi(g+1)=toList(\Phi(g))$
           - Add $N-|\Phi(g)|$ additional parameterizations to $\Pi(g+1)$ using genetic evolution on $\Pi(g)$
     (4) $g=g+1$
  end while
  $\Pi(G)=$ Resample $\Pi(G-1)$ proportional to fitness
  return $\Pi(G)$
EPDM learns this distribution of parameterizations using genetic evolution. Initially, a set of $N$ parameterizations (denoted $\Pi(0)$) are randomly generated. Over a series of $G$ generations, the pool of parameterizations is evolved to better estimate the distribution of strategies used by human players in the supplied player games. Each generation $g$ proceeds in three steps. In step 1, the set of player games for which parameterization $\pi\in\Pi(g)$ is a top performer is computed. This set, denote $\Gamma^{\pi}(g)$, is determined as follows. Let $a^{\gamma}(t)$ denote the token allocations made in player game $\gamma\in\Gamma$ in round $t$. Also, let $a^{\pi}(t)$ be the allocations that parameterization $\pi$ would make in round $t$ given the history of game transactions made prior to round $t$. Then, the average error of $\pi$ in player game $\gamma$ is
$\delta^{\pi}_{\gamma}=\frac{1}{T_{\gamma}}\sum_{t=1}^{T_{\gamma}}dist(a^{\gamma}(t),a^{\pi}(t)),$ (1)
where $T_{\gamma}$ is the number of rounds in player game $\gamma$ and $dist(a,b)$ is the error function specifying the distance between allocations $a$ and $b$ (see SM-D for details). $\pi$ is a top-performing parameterizations for player game $\gamma$ if $\delta^{\pi}_{\gamma}-\min_{\pi^{\prime}\in\Pi}\delta^{\pi^{\prime}}_{\gamma}<\varepsilon$. We used $\varepsilon=0.01$.
In step 2 of Algorithm 1, a core set of parameterizations $\Phi(g)\in\Pi(g)$ is computed using the greedy process defined in the function FindCoreSet (see SM-C.2). FindCoreSet repeatedly adds the parameterization in $\Pi(g)$ that is a top-performer in the most player games that have not yet been represented until all player games are represented by at least one top-performing parameterization.
Finally, in step 3, the EPDM algorithm constructs the set of parameterizations used in the subsequent generation. First, each parameterization from the core set $\Phi(g)$ is added to $\Pi(g+1)$. The remaining parameterizations then are generated using genetic evolution on the parameterizations in $\Pi(g)$. That is, parameterizations are evolved using selection, crossover, and mutation (see SM-C.3 for details). In this process, the fitness of parameterization $\pi$ is the cardinality of $\Gamma^{\pi}(g)$ (i.e., $|\Gamma^{\pi}(g)|$.
After all generations are completed, the final distribution of parameterizations returned by the EPDM algorithm is determined by resampling $\Pi(G-1)$ based on fitness. See SM-C.4 for details.
4 Experiment A: Evaluating population dynamics
Our goal is to find models of human behavior in the JHG that satisfy two criteria. First, the population dynamics of agent collectives should mirror the population dynamics of human societies. Second, individual agents should display human-like behavior. In this section, we evaluate four models, created by combining parameter-based models (hTFT and hCAB111hCAB and hTFT denote parameterizations of CAB and TFT models, respectively, learned from human data.) with modeling algorithms (PSO and EPDM), with respect to the first criteria. In Section 5, we address the second criteria.
4.1 Experiment design
We simulated societies of hTFT-PSO, hTFT-EPDM, hCAB-PSO, and hCAB-EPDM agents in the JHG, and compare them to human societies. We also consider two baseline agents: (1) Random, which randomly allocates tokens with percentages of give, take, and keep that match the percentages used in the training set (Section 4.2) and (2) eCAB agents [41]. eCABs are CAB agents with parameterizations learned via evolutionary simulations rather than from human data. eCABs display somewhat effective behavior in the JHG, though their strategies are distinct from humans [41].
To evaluate these agents, we compare their population dynamics with those of human players observed in the 15 games in our test set (Section 4.2). Simulations for agent populations were conducted with identical settings (initial popularities, number of agents, and game length) as games found in the test set. For each game in the test set, four simulations of each agent population were conducted with identical numbers of agents and rounds, bringing the total number of simulations for each agent to 60. In setting up experiments in this way, we can directly compare human and agent societies.
For simulations with hTFT-EPDM agents, hCAB-EPDM agents, and eCAB agents, parameterizations were randomly selected from the learned pool of strategies for each game. On the other hand, for hTFT-PSO and hCAB-PSO, the parameterizations to which the PSO algorithm converged from four runs of the PSO algorithm were used.
4.2 Data sets
Our data set, which is provided in the supplementary material, consists of 66 games played primarily by experienced human players. 47 of these 66 games had only human players. The other 19 games were played by both human and bot players. Of the 47 all-human games, we randomly selected 15 of these games for the test set. The remaining 51 games were designated as training data. Games in the training set had 5-12 players with an average of 8.08 players (of which 6.57 were humans). Across these games, we have a total of 335 human player games (which comprise $\Gamma$). The average number of rounds in these player games was 24.96. Games in the test set were played by 6-11 human players (average was 7.93 players). Games with more than 30 rounds were truncated at 30 rounds.
4.3 Metrics to evaluate population dynamics
Table 1: Metrics used to compare the population dynamics of human and agent societies.
Category Description Metrics
Wealth & Power Wealth and influence of society members Mean popularity, Gini Index
Economic The high-level actions (give, take, keep) % give, % take, % keep,
Behavior taken by individuals in society Evolution Coefficient
Mixing Connections across society, including Reciprocity, Density,
Patterns dyadic, triadic, and group relationships Entropy, Polarization
Summary Comparison to humans across all metrics Mahalanobis distance
We evaluate population dynamics using the metrics summarized in Table 1. These metrics are grouped into four categories, each designed to evaluate a particular aspect of population dynamics. Metrics in the first category quantify the wealth and power wielded by society members. We use Mean Popularity, an approximation of Katz centrality [24], and Gini Index, which measures the distribution of wealth across society. Economic behavior summarizes player actions across society, including the percentages of token allocations of each type (keep, take, and give) and the average change in playar token allocations from round to round (Evolution Coefficient). Next, we assess societal mixing patterns, or the way that players both join together in pairs and groups, and fragments and segregates. These mixing patterns are quantified using two measures of reciprocity, density, entropy, and polarization. Finally, we summarize the dissimilarity between agent and human populations across all these metrics using Mahalanobis distance [27]. Formal definitions of each metric, as well as motivation for why we chose these (as opposed to other) metrics, are provided in SM-E.
4.4 Results
Figure 2 displays measures of human and agent population dynamics. We discuss each by category.
(a)                                     (b)                                     (c)
Society % Give % Keep % Take
Humans   75 19   6
Random   72 24   4
eCAB   95   5   0
hTFT-PSO   97   3   0
hTFT-EPDM   70 15 15
hCAB-PSO 100   0   0
hCAB-EPDM   74 19   7
(d)                                   (e)                                       (f)
Figure 2: Comparisons of human and agent societies. Dots (a-c), numbers (d), and lines (e-f) denote the mean, error ovals (which do not account for covariance) and ribbons show the standard error.
Wealth and power: Figure 2a shows a comparison of agent and human populations with respect to Mean Popularity and Gini Index. In these two regards, hCAB-EPDM populations most closely mirror those of humans. Random populations tend to have both low Gini Index and low mean popularity. hTFT-EPDM populations tend to have slightly higher Gini Index and lower popularity, while hTFT-PSO, hCAB-PSO, and eCAB populations all have higher mean popularity and lower Gini Index. While high average popularity and lower Gini Index is desirable (though this does not necessarily indicate robust behavior in general), our objective is to identify models that mimic human behavior. With respect to mean popularity and Gini Index, the CAB model coupled with EPDM parameter modeling results in the agent model that most closely reflects human population dynamics.
Economic behavior: In the JHG, wealth outcomes are the result of economic behavior. eCAB agents and algorithms that model parameters using PSO all have higher mean popularity and lower Gini Index than human populations. These agents use most of their tokens to give to each other. They keep and take little, if any (Figure 2d). On the other hand, humans and agents modeled with EPDM have higher variability in strategies across players and tend to keep and take more tokens. This produces greater inequality in the society and less overall prosperity. As such, agents produced using EPDM tend to make more human-like distributions of token allocations.
The evolution coefficient measures the consistency of player token allocations over rounds. Figure 2e shows that none of the agent populations display similar behavior to humans in this regard. CAB agents are too consistent in their allocations (though hCABs are less consistent than eCABs), showing less change in token allocations from round to round than humans. On the other hand, TFT agents display a sort of multiple personality disorder. The act of reciprocating actions from the previous round produces oscillating changes in token allocations that are clearly not human-like. However, the differences in token allocations of TFT agents over two rounds are quite small.
Mixing patterns: We measure two forms of reciprocity: Immediate Reciprocity (reciprocation from one round to the next) and Overall Reciprocity (reciprocation over a complete game). In the JHG, humans have moderate levels of both forms of reciprocity (Figure 2b). EPDM agents exhibit similar levels of reciprocity, with hCAB-EPDM’s rates being only slightly higher than people’s. On the other hand, eCABs and PSO reciprocation rates are far higher than those of humans (in the data set).
The density and entropy measures, shown in Figure 2c, provide information about the number (density) and strength (entropy–are all trading partners valued similarly?) of trading partners. In these regards, hCAB-EPDM agents once again tend to be very similar to those of human players. hCAB-PSO agents have similar density, but higher entropy. On the other hand, agents using TFT tend to be less similar to humans in both regards. Thus, hCAB agents more effectively mirror these human behaviors better than TFT agents, with hCAB-EPDM being the best.
Finally, we consider polarization (Figure 2f), which describes the degree to which players segregate into groups. Since the CAB algorithm is based on group formation, while TFT is not), it is not unexpected to observe that all CAB agents display higher polarization than TFT agents. These higher levels are similar to the polarization measures we observe in human games. This again shows that the CAB model can be tuned to better model human dynamics than the TFT (matching) model.
Metric Random eCAB hTFT-PSO hTFT-EPDM hCAB-PSO hCAB-EPDM
Mahalanobis Dist. 18.822   7.089 11.565   7.656   9.216   3.393
Chi-Square Test $p<0.001$ $p<0.001$ $p<0.001$ $p<0.001$ $p<0.001$ $\mathbf{p=0.486}$
Table 2: Mahalanobis distance (lower is better) between human and agent populations over all metrics. $p>0.05$ indicates that no difference is detected between human and agent populations.
Summary metric: Table 2 gives the Mahalanobis distance between the various agent populations and our human population across the eleven individual metrics. Of the six agent populations, hCAB-EPDM populations most closely mirror human societies from our test set. Indeed, a Chi-Square test detects no statistical difference between hCAB-EPDM and human populations ($p=0.486$), thus indicating a relatively close fit across these measures. All other agent populations were statistically distinct from human populations ($p<0.001$ in all cases).
These results show that the CAB algorithm paired with EPDM parameterization modeling produces agent populations whose dynamics are surprisingly close (with some variation–notably the evolution coefficient) to those of human populations in the JHG (first criteria). In the next section, we consider whether individual hCAB-EPDM agents plausibly exhibit human behavior (second criteria).
5 Experiment B: Interacting with human players
In this section, we seek to establish the degree to which individual hCABs behaviors seem human-like to people. To do this, we conducted a user study pairing human players with hCABs in the JHG.
5.1 User study design
We recruited eight people, each of which had prior experience playing the JHG, to participate in the study. The study proceeded in a series of four periods. In each round, the participants were randomly divided in half. Each group of four participants was paired with four (randomly selected) hCABs to form two games played by eight players each. Thus, eight games were played across the four periods. Games lasted for 21, 20, 17, and 15 rounds in Periods 1-4, respectively (players were not told game lengths). For consistency, only the first 15 rounds of each game were used in our data analysis.
Participants were asked to play so as to become as popular as possible by the end of each game, while also identifying whether associates were humans or bots. Participants were told in advance that each game was composed of four people and four bots. After each game, participants completed a survey asking them to identify each associate as a human or a bot. We also compare humans and hCABs with respect to their popularity and the percentage of tokens used to give, keep, and take.
5.2 Results
Society % Give % Keep % Take
Humans   60 25 15
hCABs   75 20   5
(b) Action percentages
                                                                      Guesser % Correct Designation Humans 56.7 Chance 57.1
                  (a) Popularity over time                           (c) Detecting humans vs. hCABs
Figure 3: User study results comparing humans and hCABs. (a) Mean popularity over time. Error ribbons show the standard error. (b) Percent of tokens used to give, keep, and take by each player type. (c) Percentage of players correctly identified by human participants verses random chance.
Over the eight games, hCABs had slightly higher average popularity than humans in rounds 5-15 (Figure 3a). This was potentially caused by some human participants being quite aggressive, attacking substantially more than hCABs (Figure 3b). Attacks sometimes lead to retaliation and prolonged fighting, resulting in lost popularity. While the distribution of human actions was different than humans in our data sets (see previous section), average hCABs behavior remained about the same.
Despite these differences between these particular human participants and hCABs, participants were unable to distinguish each other from hCABs on average. Figure 3c shows that participants correctly identified whether their associates were human or bot 56.7% of the time. Given that participants knew that four of their associates were bots and three were humans, random chance would have produced a correct detection rate of 57.1%. Therefore, participants were no better than random in identifying whether there associates were humans or hCABs. This gives credence to the argument that the behaviors of the randomly selected hCABs seem human-like.
Viewing hCAB behavior in individual games provides further evidence of this claim. For example, consider Figure 4, which depicts one of the eight games from the study (similar graphs for the other games are provided in SM-G). In Round 9, two human players (p1 and p2) attacked each other. Interestingly, three of the four hCABs and the two other humans all joined with p7 in attacking p1 in Round 10 (which greatly weakened p7 for the remainder of the game). That the hCABs all chose to take the same side in the dispute between p7 and p1 as the other human players (siding with the better connected individual) illustrates the ability of hCABs to reason about groups and power in a way that led them to mirror human behavior in this situation. Similar observations can be made about how hCABs joined together in subgroups with humans and other hCABs in these games.
Figure 4: Illustrative game showing the popularity of players over time (top) and network structure in select rounds (bottom). In the graphs, arrows indicate token transactions in the current round (green = give; red = take), while nodes more connected historically tend to be closer to each other.
6 Conclusions and delimitations
In this paper, we compared and contrasted several different algorithms in modeling human behavior in a strategic network game called the Junior High Game (JHG) from a relatively small data set. We considered approaches that differ with respect to the underlying assumptions of human behavior they use (behavior matching verses community-aware behavior) and the statistical moments they model (mean verses distribution). Our results indicate that a parameter-based community-aware model (CAB [41]) coupled with evolutionary population distribution modeling (EPDM) can produce agents that closely reflect human population dynamics in the JHG. Furthermore, results of a user study indicate that the individual behaviors of these agents plausibly mimic human decision-making.
The CAB model consists of a variety of hand-coded reasoning functions modulated by thirty parameters. Such a modeling approach has advantages and disadvantages compared to a more tabula rasa approach (such as a graph neural network). One advantage is that an effective model of human behavior can be derived using only a small data set. This is valuable, since obtaining large amounts of quality data in strategic network games can be challenging. A second advantage is that the trained strategies are interpretable, thus potentially shedding insight on human behavior. On the other hand, a tabula rasa approach could potentially learn other kinds of strategies, and is therefore a potentially promising area for future work. hCAB-EPDM provides a baseline for such future methods.
While our results indicate that hCAB-EPDM agents are, perhaps surprisingly, effective models of human behavior in the JHG, we caution against overly broad extrapolations. For example, the societies modeled in this paper were small (6-11 players). Larger societies have different characteristics. Additionally, the hCAB-EPDM agents used in this paper were trained on a small data set. Given the complex state and strategy space of the JHG, other human players could potentially use different strategies not found in this data set. Though the CAB algorithm can be configured (via its parameters) to model many different kinds of behavior, it is possible that other communities of human players could display vastly different norms of acceptable behavior that may not be effectively modeled by parameterizations of the CAB algorithm. Future work could consider whether behavior in larger, more varied, data sets could be modeled effectively by CAB agents. Finally, the value of the results described in this paper somewhat hinges on the degree to which the strategic network game we selected (the JHG) models human networks. While the JHG models important properties of such networks, the world is certainly more complex than any game can model. Future work should consider refinements and enhancements to study other complexities of human networks. These and other future works can potentially help us continue to better understand and model human networks.
Appendix A Overview of the Supplementary Material
The supplementary material consists of two parts:
  1. 1.
    Code and data sets used on the study. These are supplied in a zip file.
  2. 2.
    The additional explanations and results provided below that support the explanations and results contained in the main paper.
The additional explanations that follow cover six topics:
  1. 1.
    Section B: Description of the parameter-based TFT algorithm.
  2. 2.
    Section C: Additional information about the EPDM algorithm.
  3. 3.
    Section D: Additional information about the error function used by the EPDM and PSO algorithms.
  4. 4.
    Section E: A formal description of the metrics used to evaluate the population dynamics of human and agent populations.
  5. 5.
    Section F: Additional notes about the first experiment reported in the main paper.
  6. 6.
    Section G: Additional notes and results related to the second experiment (user study) reported in the main paper.
Appendix B A parameterized-based tit-for-tat agent for the JHG
In this section, we defined the parameter-based matching algorithm (tit-for-tat; TFT) discuss in this paper to model human behavior in the JHG.
Reciprocating the behavior of associates in the JHG is not as straightforward as it is in repeated games. A TFT algorithm for the JHG must specify three things. First, the algorithm must specify how the player will allocate tokens in the first round. Second, the algorithm must determine what the player will reciprocate. A player could reciprocate based on either the number of tokens given (or taken) or the amount of influence (number of tokens multiplied by popularity) given (or taken). Third, the algorithm must determine what to do when agents receive more or less tokens from other players than they are able to reciprocate.
The seven parameters specified in Table 3 dictate the token allocations made by the agent. In the hTFT agents used in this paper (hTFT-PSO and hTFT-EPDM), these parameters are tuned to match behavior specified in the training set as closely as possible using the PSO and EPDM modeling methods, respectively.
We define how TFT agents act given these parameters in this section. The algorithm is implemented in the file TFTAgent.h in the supplied code.
Table 3: The set of parameters $\Theta$ used in the TFT parameter-based algorithm. Each parameter can take on an integer value in the range $[0,100]$.
Symbol Usage
$\theta_{\rm initKeep}$ Specifies the percentage of its tokens the agent keeps in the first round
$\theta_{\rm initAllocSize}$ Specifies the average number of tokens allocated to selected associates in the
first round
$\theta_{\rm initPercNeg}$ Specifies the probability that an initial token allocation is positive (give) or
negative (attack)
$\theta_{\rm matchType}$ Determines whether the agent matches tokens ($<50$) or influence ($\geq 50$)
$\theta_{\rm notEnough}$ Determines, in the case of not being able to reciprocate everything, whether the
agent scales back all token allocations proportionally ($<50$) or prioritizes higher
impact relationships ($\geq 50$)
$\theta_{\rm tooMany}$ Determines what to do in case of having excess tokens. When $<50$, the agent
scales up all token allocations proportionally. Otherwise, it creates new random
allocations.
$\theta_{\rm keepExtra}$ Specifies the percentage of extra tokens (after reciprocating everything) that the
agent keeps
B.1 First-round token allocations
The first three parameters in Table 3 ($\theta_{\rm initKeep}$, $\theta_{\rm initAllocSize}$, and $\theta_{\rm initPercNeg}$) impact how the agent allocates tokens in the first round. The following rules define the agent’s behavior in this round:
  • $\theta_{\rm initKeep}$ specifies the percentage of tokens the agent keeps in the first round. That is, the agent keeps $\left(\frac{N\cdot\theta_{\rm initKeep}}{100}\right)$ tokens, rounded to the nearest integer, in the first round. Here, $N$ is the number of tokens it has to allocate. We used $N=2|I|$ (two times the number of players) in our games.
  • The remaining tokens are allocated by randomly selecting players in the game (who will be the recipient of an allocation) and then allocating a subset of tokens to them. The number of tokens allocated to the selected agent is $z=1+{\rm round}\left((N-1)*\frac{R}{100}\right)$, where ${\rm round}(\cdot)$ rounds the value to the nearest integer, $R=\max(0,\min(100,\theta_{\rm initAllocSize}+r_{20}-10$)), and $r_{20}$ is an integer randomly selected in the interval $[0,20]$. If $z$ is greater than the number of tokens the agent has remaining to allocate, it is truncated to the remaining number of tokens.
  • The allocations made in the previous bullet point can be positive (give) or negative (attack). $\theta_{\rm initPercNeg}$ specifies the probability that the agent makes each allocation positive or negative.
B.2 Determining what to reciprocate
In the JHG, there are two ways one can think about matching behavior. First, the agent can just reciprocate the number of tokens (positive or negative) that a player allocated to them in the previous round. However, because the tokens of more popular players are more influential than the token allocations of less popularity people, an agent could also choose to match influence instead of number of tokens. The influence exchanged in a token allocation is give by the popularity of the allocator $P_{i}$ multiplied by the number of tokens allocated. $\theta_{\rm matchType}$ specifies whether the TFT agent matches tokens or influence. If $\theta_{\rm matchType}<50$, then the agent matches tokens. Otherwise, it matches influence.
B.3 Dealing with excess and shortfall
Regardless of what the agent chooses to reciprocate (tokens or influence), it can receive more or less in a round than it has the ability to reciprocate. We deal with each case individually.
First, in the case of receiving more than it can reciprocate, the agent uses one of two methods. When $\theta_{\rm notEnough}<50$ the agent simply computes the number of tokens needed to match all transactions from each agent, and then scales back all allocations proportionally. When $\theta_{\rm notEnough}\geq 50$, the agent first reciprocates fully with agents with whom it has historically had the most interactions, determined by $|{\mathcal{I}}(i,j)|+|{\mathcal{I}}(j,i)|$, where $|{\mathcal{I}}(i,j)|$ is the influence of player $i$ on player $j$ in the JHG influence matrix [41]. It continues to fully reciprocate to those agents until it has no remaining tokens. Thus, it will not tend to reciprocate the actions of associates with whom the overall strength of its interactions have been small in the past.
Second, in the case of having more tokens than are needed to reciprocate all of last round’s transactions, the agent first ensures that it reciprocates all allocations made to it in the last round by other players. It then keeps a percentage of the remaining tokens specified by $\theta_{\rm keepExtra}$. Finally, with the $100-\theta_{\rm keepExtra}$ percent of the remaining tokens, it simply scales up all token allocations proportionally when $\theta_{\rm tooMany}<50$, or, otherwise, randomly allocates tokens (positively or negatively based on $\theta_{\rm initPercNeg}$) to randomly selected associates.
Appendix C EPDM
C.1 Implementation
The implementation of the EPDM algorithm we used is given in the file EPDM.cpp of the supplied code. Note that we used a gene pool of size $N=100$ parameterizations in each generation.
C.2 FindCoreSet function
Due to space limitations, the FindCoreSet function was not provided in the main paper. This algorithm is given in Algorithm 2 of this supplementary material. Variables are defined as in the main paper.
Algorithm 2 FindCoreSet – Selects a set of parameterizations from the current set of parameterizations $\Pi(g)$ to describe the set of observed human strategies.
  Input:
     - $\Gamma$ (set of player games)
     - $\Pi(g)$ (list of parameterizations used in generation $g$)
     - $\Gamma^{\pi}(g)$ for each $\pi\in\Pi(g)$ (player games for which $\pi$ is a top-performing parameterization)
  Initialize:
     - $\hat{\Gamma}=\Gamma$
     - $\hat{\Gamma}^{\pi}(g)=\Gamma^{\pi}(g)$
     - $\Phi=\emptyset$
  while $\hat{\Gamma}\neq\emptyset$ do
     $\pi^{*}=\max_{\pi\in\Pi(g)\setminus\Phi}|\hat{\Gamma}^{\pi}(g)|$
     $\hat{\Gamma}=\hat{\Gamma}\setminus\hat{\Gamma}^{\pi^{*}}$
     $\Phi=\Phi\cup\{\pi^{*}\}$
     $\forall\pi\in\Pi(g),\hat{\Gamma}^{\pi}(g)=\hat{\Gamma}^{\pi}(g)\setminus\hat{\Gamma}^{\pi^{*}}(g)$
  end while
  return $\Phi$
C.3 Details of evolutionary algorithm
Step 3 of the EPDM algorithm (see Algorithm 1 in the main paper) specifies that new parameterizations are added to $\Pi(g+1)$ (the parameterization/gene pool for the next generation) using genetic evolution. As in standard of genetic evolution, this algorithm uses selection, crossover, and mutation. This was done as follows:
  1. 1.
    Two parameterizations are randomly (based on fitness) selected from $\Pi(g)$. The probability a parameterization $\pi$ is selected is given by $\frac{|\Gamma^{\pi}(g)|}{\sum_{\pi^{\prime}}|\Gamma^{\pi^{\prime}}(g)|}$, where $|\Gamma^{\pi}(g)|$ is the cardinality of the set $\Gamma^{\pi}(g)$ (the number of player games for which $\pi$ is a top performer).
  2. 2.
    A new parameterization is formed from the two parameterizations selected in the prior step. In this parameterization, the $i^{\rm th}$ value is determined in one of three ways. With probability 0.03, it is a random value in the range $[0,100]$. Otherwise, the value is chosen to be the $i^{\rm th}$ value of one of the two selected parameterizations (chosen randomly). With probability 0.12, this value is then randomly shifted up or down from $[0,5]$ values (truncated to be between 0 and 100, inclusive).
C.4 Resampling
The last step of the EPDM algorithm (Algorithm 1 from the main paper) consists of resampling the parameterizations evaluated in the last generation $\Pi(G-1)$ based on fitness. This is done by randomly selecting $N$ parameterizations from the list $\Pi(G-1)$. In selecting each parameterization, the probability that parameterization $\pi$ is chosen is given by $\frac{|\Gamma^{\pi}(g)|}{\sum_{\pi^{\prime}}|\Gamma^{\pi^{\prime}}(g)|}$. Note that a parameterization can be chosen more than once.
Appendix D Error Functions
Both the PSO and EPDM algorithms for learning parameterizations from data require an error or distance function (e.g., Eq. 1 in the main paper). This function takes as input two token allocation profiles, and returns a value specifying the difference between them. We experimented with several such functions, including mean-squared error (MSE) and a custom built function, called property scoring (PS). In PS, the error is based on how well the two token allocation profiles match across a variety of different properties. More specifically, the PS error function scores pairs of token allocation profiles using a set of properties. The distance between the actual allocations and the proposed allocations is decreased for keeping the right number of tokens, giving the right number of tokens in total, giving tokens to the right players, giving the right number of tokens to the right players, taking at the right times, taking the right number of tokens, and taking from the right player. The distance function is increased for proposing to give to a player from which tokens were actually taken or taking tokens from a player for which tokens were actually given.
Exact specifications of the property scoring distance function are given in the file EPDM.cpp (see the function scoreProposedAllocation) of the supplied code.
Because the PS error function always performed better than MSE, we used this scoring function in the training of all agent models.
Appendix E Metrics
In this section, we define how we computed each of the metrics used in the paper to evaluate population dynamics. These metrics are derived from the known quantities of the JHG, including the transaction matrix $X$, the popularity vector ${\mathcal{P}}$, and the influence matrix $\mathcal{I}$. Additionally, to capture longer-term (non-discounted) interactions between players, we use the impact matrix. We define the impact matrix prior to discuss each of the metrics used in the paper.
We define the economic actions taken by a player at time $\tau$ as a matrix $X(\tau)$. Some metrics, discussed below, use both positive and negative connections while others only one or the other. To distinguish between each case while trying to simplify explanations we distinguish between $X(\tau)$, $X^{+}(\tau)$, and $X^{-}(\tau)$ as all connections, positive connections, and negative connections. We apply this to the different derived matrices as well, such as the influence matrix or impact matrix presented below.
E.1 The impact matrix
To examine the evolution of friendships and interaction patterns throughout the Junior High Game, we construct the impact graph, a network representation that aggregates player interactions over some duration of the game. This graph encodes social mixing patterns by weighting interactions based on their frequency, the relative popularity of the involved players at the time of interaction, and the corresponding economic action coefficient.
To calculate the impact matrix, we first calculate $J^{\prime}_{ij}(\tau)$ representing the interaction strength between players $i$ and $j$ at time step $\tau$. It is defined as:
$J_{ij}(\tau)=\begin{cases}\mathcal{P}(\tau)X_{ij}(\tau)c^{keep},&\text{if }i=j\\ \mathcal{P}(\tau)X_{ij}(\tau)c^{give},&\text{else if }x_{ij}\geq 0\\ \mathcal{P}(\tau)X_{ij}(\tau)c^{take},&\text{else if }x_{ij}<0\end{cases}$ (2)
where $\mathcal{P}_{i}(\tau)$ represents the popularity of player $i$ at time $\tau$ adjusting the weight of interactions based on social influence. $X_{ij}(\tau)$ is the economic action between players $i$ and $j$ with $c^{give}$, $c^{keep}$, and $c^{take}$ representing the economic action coefficients that assign weights based on the type of interaction (giving, keeping, or taking) respectively.
The aggregated impact matrix over a time interval $[q,r]$ is given by:
$J^{\prime}(q,r)=\sum_{\tau=q}^{r}J(\tau)$ (3)
which sums the impact matrices over all time steps within the interval. We then normalize the interaction weights so that we can compare relative influences across different context as follows:
$J^{\prime}(q,r)=\frac{J(q,r)}{\|J(q,r)\|_{1}}$ (4)
where $\|J^{\prime}(q,r)\|_{1}$ is the L1 norm of the matrix, ensuring that the sum of the magnitude of all entries in $J(q,r)$ is constrained within a unit scale.
Additionally, we often look at the impact of just one round. That is, if $q=r$, we use the notation $J^{\prime}(r)$ to mean $J^{\prime}(q,r)$ to simplify understanding of our equations.
E.2 Metrics for wealth and power
Because the JHG uses popularity (an approximation of Katz centrality [24]) as the primary commodity, we utilize the mean popularity, $\mu_{\mathcal{P}}$, as a metric to quantify the productivity of a society. To apply this measure, we first constructed a popularity vector $\mathcal{P}=[\mathcal{P}_{0},\mathcal{P}_{1},\ldots,\mathcal{P}_{n-1}]$ representing the popularity of each agent $i$ at a particular time $\tau$. We then define average popularity as follows:
$\mu_{\mathcal{P}}(\tau)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{P}_{i}(\tau)$ (5)
We additionally employ the Gini Index, $G(\tau)$, as a quantitative metric to assess inequality of popularity across a society. The Gini Index, originally developed as a measure of income inequality [16], ranges from 0 to 1, where 0 indicates perfect equality (i.e., all individuals possess equal popularity) and 1 denotes maximal inequality (i.e., one individual possesses all of the popularity while others have none). $G(\tau)$ was then computed using the following standard formula:
$G(\tau)=\frac{1}{2n^{2}\mu_{\mathcal{P}}(\tau)}\sum_{i,j=0}^{n-1}|\mathcal{P}_{i}(\tau)-\mathcal{P}_{j}(\tau)|$ (6)
where $n$ is the number of agents and $\mu_{\mathcal{P}}$ is the mean of the values in vector $\mathcal{P}$. In our analysis, we used the Gini Index (as opposed to the variance of the population) because it is a preferred measure used in measuring inequality in societies in economics [38, 2] and sociology [28].
We used average popularity and the Gini Index as an interpretable and sensitive measures for assessing the mean and spread of popularity in a society.
E.3 Metrics for quantifying of economic behavior
To help understand how individuals in a society interact with each other, we analyze their economic actions using the percent of give, keep and take allocations over societies of each type of agent. Precise definitions of these metrics are as follows:
[…]
where $X_{ij}(\tau)$ represents the token allocations of player $i$ toward player $j$ at time $\tau$. The diagonal of the transaction matrix $X(\tau)$ specifies agent keeping, while positive non-diagonal values represent giving and negative non-diagonal values represent taking. $T$ is the number of rounds in the game, $|I|$ is the cardinality of the set $I$ (specifying the number of players in the game), and $\delta$ is the Kronecker delta function.
E.3.1 The evolution coefficient
To capture the evolution of economic behaviors in the JHG over time, we use the evolution coefficient. The evolution coefficient quantifies structural changes in the impact matrix over time. Given that the impact matrix $J^{\prime}(q,r)$ represents normalized interaction patterns over a specified window. This coefficient, denoted $C_{d}$, is computed by evaluating differences in consecutive impact matrices, shifted by a delay $d$:
$C_{d}=\sum_{\tau=0}^{r-p-1}\sum_{ij}\left|J^{\prime}_{ij}(\tau)-J^{\prime}_{ij}(\tau+d)\right|$ (10)
In words, the evolution coefficient quantifies the cumulative absolute differences between the impact matrices over a specified number of rounds $d$. A higher value of $C_{d}$ indicates more substantial shifts in the interaction structure, suggesting dynamic social mixing or evolving economic strategies among players. Conversely, lower values suggest stability in social interactions and player behaviors.
We note that, in the computation of the Mahalanobis distance, the evolution coefficient is measured as the average change in token allocations across one and two rounds of play.
E.4 Mixing patterns: Measuring reciprocation
Reciprocity is an important component of cooperative dynamics within agent societies [31]. High reciprocity suggests that players consistently return favors and maintain balanced exchanges, whereas low reciprocity indicates asymmetric relationships, where some players give or take more than they receive. We define the pairwise global reciprocity between agents $i$ and $j$ over the course of an interaction as:
$R^{global}_{ij}=\begin{cases}0&\text{if }i=j;J_{ij}(r_{0},r_{f})\leq 0;J_{ji}(r_{0},r_{f})\leq 0\\ J^{\prime}_{ji}(r_{0},r_{f})\frac{J_{ij}(r_{0},r_{f})}{J_{ji}(r_{0},r_{f})}&\text{else if }J_{ij}(r_{0},r_{f})<J_{ji}(r_{0},r_{f})\\ J^{\prime}_{ji}(r_{0},r_{f})&\text{else if }J_{ij}(r_{0},r_{f})\geq J_{ji}(r_{0},r_{f})\\ \end{cases}$ (11)
Here, $J_{ij}(r_{0},r_{f})$ denotes the cumulative reciprocated impact from agent $j$ to agent $i$ over the interval $[r_{0},r_{f}]$, where $r_{0}$ if the first round of the game and $r_{f}$ is the final round of the game. The metric $R^{global}_{ij}$ thus captures the extent to which agent $j$ reciprocates the contributions made by agent $i$, normalized by the directionally lesser interaction.
We use the overall global reciprocity coefficient to define overall reciprocity. It is summarizes the aggregate reciprocity across all agent pairs in the population as follows:
$R^{global}_{coeff}=\frac{\sum_{ij}{R^{global}_{ij}}}{n(n-1)}$ (12)
where $n=|I|$ denotes the number of players in the game. This coefficient provides a population-level estimate of the prevalence of mutual exchanges, weighted by the magnitude of impact transferred between agents over the duration of the game.
In addition to this cumulative measure, we also define immediate reciprocity to capture short-term reciprocation dynamics. The pairwise local reciprocity at round $r_{m}$ is defined as:
$R^{local}_{ij}=\begin{cases}0&\text{if }i=j;J_{ij}(r_{m})\leq 0;J_{ji}(r_{m})\leq 0\\ J^{\prime}_{ji}(r_{m-1})\frac{J_{ij}(r_{m})}{J_{ji}(r_{m-1})}&\text{else if }J_{ij}(r_{m})<J_{ji}(r_{m+1})\\ J^{\prime}_{ji}(r_{m-1})&\text{else if }J_{ij}(r_{m})\geq J_{ji}(r_{m-1})\\ \end{cases}$ (13)
The corresponding local reciprocity coefficient is:
$R^{local}_{coeff}=\frac{\sum_{ij}{R^{local}_{ij}}}{n(n-1)}$ (14)
This formulation evaluates how frequently short-term exchanges are reciprocated from one round to the next, capturing the temporal responsiveness of agents to recent interactions.
E.5 Mixing patterns: Measuring connectivity using density and entropy
Graph density provides a measure of how interconnected a network is at a given time, relative to the maximum possible number of directed edges (excluding self-loops). Just to note, when average across the whole network average in/out degree are proportional to density and therefore we exclude the redundant comparison. Density is defined as:
$D(\tau)=\frac{1}{n(n-1)}\sum_{i,j=0;i\neq j}^{n-1}A^{+}_{ij}(\tau)$ (15)
Here, $A^{+}_{ij}(\tau)$ represents an adjacency matrix which is $1$ if a positive economic action from node $i$ to node $j$ took place in round $\tau$. Otherwise it is 00. The denominator $n(n-1)$ reflects the maximum number of actions in a loop-free graph.
While density describes the average number of associates a player is interacting with, entropy quantifies the distribution of a player’s actions among agents. We define a normalized entropy score that captures how evenly distributed its outgoing edge weights are. For node $i$, let the set of positive outgoing edge weights be $\{X^{+}_{ij}\}_{j=1}^{n}$. Define the normalized edge weight distribution as:
$w_{ij}=\frac{X_{ij}^{+}}{\sum_{k=1}^{n}X_{ik}^{+}}$ (16)
The entropy for node $i$ is then given by:
$H_{i}=-\sum^{n-1}_{j=0;X^{+}_{ij}>0}w_{ij}\log w_{ij}$ (17)
The maximum possible entropy $H_{i}^{\text{max}}$ occurs when all outgoing edges from node $i$ are equally weighted, i.e.,
$H_{i}^{\text{max}}=\log k_{i},$ (18)
where $k_{i}$ is the out-degree of node $i$ (number of positive-weight outgoing edges). Then entropy is then computed as:
$H_{i}^{\text{norm}}=\begin{cases}0&\text{if }\sum_{j}X^{+}_{ij}=0\text{ or }k_{i}=1\\ \frac{H_{i}}{H_{i}^{\text{max}}}&\text{otherwise}\end{cases}$ (19)
Finally, the average entropy across all nodes provides a summary statistic for the overall spread of edge weight distributions in the network:
$H^{\text{avg}}=\frac{1}{n}\sum_{i=1}^{n}H_{i}^{\text{norm}}$ (20)
E.6 Measuring group formation: Polarization
Finally, we desire to quantify how players cluster and separate. Known metrics for (partially) quantifying such behaviors include the clustering coefficient [47, 29] and modularity [30]. However, many metrics either (1) are not designed for signed, weighted, and time-varying graphs, (2) do not provide comparisons of grouping behavior from graph to graph (e.g., modularity measures how well a community partition describes a graph, not how divided or grouped a graph is), or (3) conflate multiple concepts (e.g., high density tends to produce a high clustering coefficient and thus may not indicate how members of society actually group and segregate). Thus, in this paper, we use a customized metric called polarization to quantify how individuals separate into different groups and sub-groups. Related somewhat to the idea of modularity, polarization measures how much individuals would consistently prefer individuals within a designated group as opposed to individuals outside of that group.
Polarization measures the maximum preference separation between two distinct (and sufficiently important) sub-groups in the society. Preference separation is a method for measuring how separated two distinct sub-groups within society are in a weighted and signed graph. Specifically, it measures how much individuals in a sub-group would prefer (based on immediate connections) any one member of their own sub-group to any one member of the other sub-group. Preference separation between two groups is highest when (1) nodes with each sub-group are well connected, (2) connections between the two sub-groups are weaker than connections within the sub-groups, and (3) the sub-groups cover a larger proportion of the entire society.
Before giving details for how preference separation and polarization are computed, we first define several quantities. Let
$m_{i}=\sum_{j}|\mathcal{I}_{ji}|$ (21)
be the magnitude of influence (both positive and negative) flowing to individual $i$ from all other individuals in society. Furthermore, let
$w_{i}=\frac{m_{i}}{\sum_{j\in I}m_{j}},$ (22)
where $I$ is the set of all individuals in the society, be the total proportion of influence that is flowing to individual $i$.
Next we define a set of dueling sub-groups in society by comparing how society would partition itself if all individuals had to choose between individuals $i$ and $j$ (where $i\neq j$. Let
$G_{ij}=\{i\}\cup\{k\in I:\mathcal{I}_{ik}>\mathcal{I}_{jk}+\delta\},$ (23)
where $\delta=0.04m_{k}$, denote the group of individuals in society that would prefer $i$ over $j$. Similarly, $G_{ji}$ denotes the group of individuals that prefer $j$ over $i$. Finally, let $G_{ij}^{\sim}=I-(G_{ij}\cup G_{ji})$ denote the set of individuals in society that are indifferent between individuals $i$ and $j$. Finally, let $W_{ij}=\sum_{k\in G_{ij}}w_{i}$ denote the collective weight of group $G_{ij}$.
Preference separation between sub-groups $G_{ij}$ and $G_{ji}$, denoted $\mathcal{S}(G_{ij},G_{ji})$, measures the proportion of individual in $G_{ij}$ and $G_{ji}$ that would prefer to stay in the same group if the choice were between other individuals from the respective groups (other than just $i$ and $j$). Formally, it is computed as follows:
$\mathcal{S}(G_{ij},G_{ji})=\left(\frac{\sum_{k\in G_{ij}}\sum_{\ell\in G_{ji}}|G_{ij}\cap G_{k\ell}|+|G_{ji}\cap G_{\ell k}|+0.5|G_{k\ell}^{\sim}|}{|G_{ij}||G_{ji}||I|}\right)\left(\frac{{\rm min}(W_{ij},W_{ji},0.2}{0.2}\right),$ (24)
where ||$\cdot$|| ⋅ | denotes the cardinality of a set. Note that the last item in the numerator of left term gives half credit when individuals are neutral between the two groups for a particular pairing. The right component of the equation scales down preference separation if one of the groups is too small (determined as having a weight of less than 0.2).
Note that measures of separation tend to be between 0.5 and 1. As such, we base polarization on a normalized version of preferences separation, given by $\mathcal{S}^{\prime}(G_{ij},G_{ji})=\frac{\mathcal{S}(G_{ij},G_{ji})-0.5}{0.5}$.
Once preference separation is computed for all pairs of sub-groups, polarization is determined to be the max preference separation for pairs of subgroups. That is, polarization (denoted $\Gamma$) is given by $\Gamma=\max_{i,j}\mathcal{S}^{\prime}(G_{ij},G_{ji})$. The constraint to only consider subgroups of a particular size ensures that the preference separation between the groups could have a meaningful impact on society.
All of the above description omit the time $t$ – polarization is computed for each time period. The actual metric we report is the average polarization over all rounds considered. We note that this metric is not computationally efficient in this formulation, but it can be computed and utilized effectively for small societies such as those we consider in this paper.
E.7 Summary metric: Mahalanobis distance
Each of the above metrics represent important aspects of population dynamics. To further compare and contrast human and agent populations, we aggregate these metrics into a multivariate Gaussian. That is, we quantify the dissimilarity between populations with respect to these metrics using the Mahalanobis distance [27]. Unlike the Euclidean distance, which assumes independence across dimensions, the Mahalanobis distance incorporates the covariance structure of the data, making it well-suited for comparing complex, interdependent societal metrics.
Given two societal types, we compute the Mahalanobis distance between them as:
$D_{M}(\mathbf{x})=\sqrt{(\mathbf{x}-\boldsymbol{\mu})^{T}\Sigma^{-1}(\mathbf{x}-\boldsymbol{\mu})}$ (25)
where $\mathbf{x}$ is the vector of observed societal metrics for a given type, $\boldsymbol{\mu}$ is the mean vector of another societal type, and $\Sigma$ is the covariance matrix of the distribution of human behavior.
To complement the Mahalanobis distance, we compute the Chi-Squared p-value, which indicates the statistical significance of the observed differences. Since the squared Mahalanobis distance follows a Chi-Squared distribution with degrees of freedom equal to the number of dimensions, the p-value provides a probabilistic interpretation of the similarity between societal types. Lower p-values suggest significant differences between the compared distributions, while higher p-values indicate that the societal structures are statistically indistinguishable.
This methodology enables us to identify when two societies exhibit fundamentally different organization patterns and to characterize the extent of these differences.
Appendix F Experiment A
Additional implementation details and results from the first experiment reported in the main paper:
  • We use the same JHG parameter settings as was used by Skaggs et al. [41]. That is, $\alpha=0.2$, $\beta=0.5$, $c_{\rm give}=1.3$, $c_{\rm keep}=0.95$, $c_{\rm take}=1.6$. Furthermore, each agent had $2*|I|$ tokens to allocate in each round.
  • eCABs were evolved with a single set of genes (rather than the 3 sets used by Skaggs et al. [41]) and were trained for 200 generations with variable initial probabilities. Simulation results were conducted with randomly selected parameterizations from the 100 parameterizations that were in the final generation.
  • Additional precision along with standard for average economic actions in each society are provided in Table 4
  • No GPU was used in the study. No high-powered computational equipment were used (only standard computing equipment with multiple cores). Training using the PSO algorithm was done on a 32-core machine, which took about 3 hours per run. Training using the EPDM algorithm took about 30 minutes on a machine with 16 cores.
Agent-Society % Give % Keep % Take
Humans $74.9\pm 1.0$ $18.8\pm 0.7$ $6.2\pm 0.8$
Random $72.0\pm 0.0$ $24.0\pm 0.0$ $4.0\pm 0.0$
hTFT-PSO $97.3\pm 1.8$ $2.7\pm 1.8$ $0.0\pm 0$ 0.0
hTFT-EPDM $69.9\pm 0.3$ $14.9\pm 0.8$ $15.2\pm 0.4$
eCAB $94.7\pm 1.8$ $5.2\pm 0.0$ $0.0\pm 0$ 0.0
hCAB-PSO $100.0\pm 0.0$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
hCAB-EPDM $73.6\pm 1.0$ $19.1\pm 0.4$ $7.3\pm 0.6$
Table 4: Mean percent of each economic actions in human and agent societies. $\pm$ values show show standard error of the mean.
Appendix G Experiment B: User study details
Additional information about the second experiment (user study) reported in the main paper:
  • The user study was approved by the authors’ institutional review board (it was given exempt status). Participants gave consent to participated in the study through the approved process.
  • Participants were paid approximately \$16 (USD) per hour while participating in the study. A high-score list (based on final popularity in each game) was kept to motivate participants to perform well.
  • Games were played using the publicly available online platform published by Skaggs et al. [41] (permission was obtained from the authors of that work to use the platform for this study). This online platform can be found at juniorhighgame.com.
  • Story plots for all eight games played in the user study are provided below. In the graphs, arrows indicate token transactions in the current round (green = give; red = take), while nodes more connected historically tend to be closer to each other.

### v3

@@
 Modeling Human Behavior in a Strategic Network Game with Complex Group Dynamics
-Abstract
+Abstract.
 Human networks greatly impact important societal outcomes, including wealth and health inequality, poverty, and bullying.
 As such, understanding human networks is critical to learning how to promote favorable societal outcomes.
-As a step toward better understanding human networks, we compare and contrast several methods for learning, from a small data set, models of human behavior in a strategic network game called the Junior High Game (JHG) [41].
-These modeling methods differ with respect to the assumptions they use to parameterize human behavior (behavior vs.
+As a step toward better understanding human networks, we compare and contrast several methods for learning models of human behavior in a strategic network game called the Junior High Game (JHG) (skaggs2024fostering).
+These modeling methods differ with respect to the assumptions they use to parameterize human behavior (behavior matching vs.
 community-aware behavior) and the moments they model (mean vs.
@@
 Results show that the highest-performing method, called hCAB, models the distribution of human behavior rather than the mean and assumes humans use community-aware behavior rather than behavior matching.
-When applied to small societies (6-11 individuals), the hCAB model closely mirrors the population dynamics of human groups (with notable differences).
-Additionally, in a user study, human participants were unable to distinguish individual hCAB agents from other humans, thus illustrating that the hCAB model also produces plausible (individual) human behavior in this strategic network game.
-1 Introduction
-Human networks, collections of people linked through social connections, influence how individuals exchange resources [26], form alliances [12, 20], and navigate competitive pressures [8, 46].
-These networks, which are studied in many disciplines [43], exhibit emergent behaviors related to inequality [6, 11, 37], information flow [9] and social behaviors [40, 10, 45].
-While the study of human networks has yielded deep insights [29, 13, 5, 35], understanding the precise decision-making processes that drive human behavior within these networks remains a major challenge [49].
+When applied to small societies, the hCAB model closely mirrors the population dynamics of human groups (with notable differences).
+Additionally, in a user study, human participants had difficulty distinguishing hCAB agents from other humans, thus illustrating that the hCAB model also produces plausible (individual) behavior in this strategic network game.
+Key words and phrases:
+Complex systems, emergent behavior, multi-agent systems, human modeling, network science
+doi:
+NQKD4743
+1. Introduction
+Human networks (collections of interconnected people) influence how individuals exchange resources willer1999network, form alliances (cranmer2012toward; gulati1995social), and navigate competitive pressures (burt2005brokerage; vu2006learning).
+These networks, which are studied in many disciplines (svendsen2009handbook; levin1998ecosystems), exhibit emergent behaviors related to inequality (barabasi1999emergence; salganik2006experimental), information flow (Harush2017), and social behaviors (shoham1997emergence; christakis2010social; centola2010spread; centola2018behavior; veenstra2021social).
+While the study of human networks has yielded deep insights (newman2003structure; easley2010networks; bar2019dynamics; pagan2019game), understanding the precise decision-making processes that drive human behavior within these networks remains a major challenge (young2015evolution).
 What determines whether an individual cooperates or competes?
@@
 Addressing these questions requires computational models that capture the complexities of human decision-making in networked environments.
-In human networks, each individual makes decisions based on personal incentives, past experiences, and the social landscape.
-Sometimes individual incentives align with the collective good, like the invisible hand [42], but other times they are diametrically opposed, like the free rider problem [34].
+In human networks, individuals make decisions based on personal incentives, past experiences, and the social landscape.
+Sometimes individual incentives align with the collective good, like the invisible hand (smith1776wealth), but other times they are diametrically opposed, as in the free-rider problem (olson1965logic).
 These individual actions lead to emergent behaviors that shape societal outcomes.
-Extensive studies have explored emergent properties of human networks [22, 23, 39], investigating, for example, how randomness affects network structures [14], how power distributes among individuals [6], and how self-organization arises within different social [47] and economic settings [1, 48].
-These macroscopic perspectives on how human societies function are valuable, yet they often are insufficient for fully modeling human networks and how they evolve.
-More detailed models of human decision-making are needed to better understand human networks.
-To help address this challenge, we consider modeling human behavior in the Junior High Game (JHG) [41], a controlled game environment designed to capture the strategic trade-offs in human networks.
-In this game, players engage in repeated economic exchanges, choosing whether to give, keep, or take resources from others.
-Decisions are shaped by social factors such as dyadic relationships [17, 15] triadic (indirect) interactions [18, 33], and group considerations [31, 7, 21].
-JHG scenarios produce networks that are dynamic, directed, weighted, and signed, thus providing a rich testbed for studying human behavior in networked settings.
-A fundamental challenge in modeling human decision-making in the JHG lies in determining how to parameterize and represent behavior [36].
+Extensive studies have explored emergent properties of human networks (holland1995hidden; jensen2022complexity; shoham2008multiagent), investigating, for example, how randomness affects network structures (erdos1959random), how power distributes among individuals (barabasi1999emergence), and how self-organization arises within different social (watts1998collective) and economic settings (arthur1994inductive; young2001individual).
+These macroscopic perspectives on how human societies function are valuable, yet they are often insufficient for fully understanding human networks and how they evolve.
+More detailed models of human decision-making on networks are needed.
+To help address this challenge, we model human behavior in a strategic network game designed to capture the emergent properties of complex (networked) societies, including wealth and power distributions and mixing patterns defined by dyadic (gouldner1960norm; fehr2000fairness), triadic (granovetter1973strength; nowak2005evolution), and group dynamics (nowak2006five; bowles2011cooperative; henrich2015secret).
+These dynamics emerge due to power asymmetry, mixed-motives, resource management, and interconnectedness easley2010networks; jensen2022complexity; thurner2018introduction, and produce networks that are dynamic, directed, weighted, and signed.
+Because these dynamics are encoded within the Junior High Game (JHG) (skaggs2024fostering), we use this strategic network game to study emergent behavior.
+A fundamental challenge in modeling human decision-making in the JHG lies in determining how to parameterize and represent behavior (rahwan2019machine).
 In this paper, we consider models that make different assumptions about human decision-making.
-One approach, behavior matching [4], assumes that agents imitate the observed actions of others, while another approach, community-aware behavior [41], considers broader structural and social cues within the network.
+One approach, behavior matching (axelrod1984evolution), assumes that agents imitate the observed actions of others, while another approach, community-aware behavior (skaggs2024fostering), considers broader structural and social cues within the network.
 We also examine how modeling different statistical moments (the mean verses the distribution of behavior) impact model performance.
-Together, this produces four different algorithms for modeling human behavior in the JHG, which we compare and contrast in small-scale societies (6-11 players).
-We evaluate the ability of the resulting modeling algorithms to produce human behavior in the JHG in two ways.
+Together, this produces four different algorithms for modeling human behavior in the JHG.
+As a starting point to understanding how well these algorithms model human behavior, we compare and contrast their behavior in small-scale societies.
+[…]
+Figure 1.
+Example scenario illustrating various JHG dynamics.
+(a) Popularity of the players over time.
+(b-c) Token allocations in rounds 1-4 and 5-6, respectively.
+Red dashed arrows indicate tokens used to attack while black solid arrows indicate giving.
+We evaluate the ability of these modeling algorithms to produce human behavior in the JHG in two ways.
 First, we compare and contrast the population dynamics of agent societies with those of human groups.
-Results show that societies of one of these agents, called hCAB, have population dynamics that closely mirror (with some differences) human populations.
-Second, we conduct a user study to identify the extent to which individual hCAB agents behave like human players.
-Results of this study show that participants were unable to distinguish hCAB agents from other human players.
-These findings not only inform the development of models that can simulate human behavior in the JHG, but also offer perspectives on the mechanisms that drive the dynamics of real-world networks.
-2 A strategic network game to model human networks
-In this paper, we model human networks using the Junior High Game (JHG) [41].
-The JHG models key properties of human networks including conflicting interests, a large and flexible strategy space, high interdependence among agents, power asymmetry, and (often) non-convergence [23, 44].
-The JHG induces a dynamic (time varying), directed, weighted, and signed network.
-We paraphrase Skaggs et al.
-[41] to review this game: The JHG is played by a set of players $I$ over a series of rounds in which each player $j\in I$ seeks to become popular.
+Results show that societies of one of these agents, called hCABs, have population dynamics that mirror (with some differences) human populations.
+Second, we conduct a user study to identify the extent to which individual hCABs exhibit human-like tendencies.
+Results of this study show that human participants have difficulty distinguishing hCAB agents from other human players.
+These findings not only inform the development of models that can simulate human behavior in the JHG, but also offer perspectives on the mechanisms that drive the dynamics of human networks.
+2. A strategic network game
+Human networks are punctuated by power asymmetry, mixed motives, resource management, and interconnectedness easley2010networks; jensen2022complexity; thurner2018introduction, which (along with other properties) combine to produce emergent properties of human networks.
+These characteristics are modeled in the Junior High Game (JHG) skaggs2024fostering, a strategic game that induces a dynamic (time varying), directed, weighted, and signed network.
+We paraphrase skaggs2024fostering to review this game: The JHG is played by a set of players $I$ over a series of rounds in which each player $j\in I$ seeks to become popular.
 Initially, every player is assigned a popularity, which changes over time based on interactions, represented by token allocations, among the players.
@@
 Finally, when $j$ attacks $i$, it negatively impacts $i$’s subsequent popularity while positively impacting $j$’s.
-After all players allocate their tokens in a round, the players’ popularities are updated based on the players’ token allocations.
+After all players allocate their tokens in a round, the players’ popularities are updated based on the token allocations.
 A new round then begins.
 The amount that a token transaction impacts subsequent popularity depends on the popularity of the player allocating the token.
-For example, receiving a token from a more popular individual results in a higher increase in popularity than receiving a token from a less popular individual.
+For example, receiving a token from a more popular player results in a higher increase in popularity than receiving a token from a less popular individual.
 Thus, players with higher popularity can have higher impact than those with lower popularity.
-(a)                                            (b)                                 (c)
-Figure 1: Example scenario illustrating various JHG dynamics.
-(a) Popularity of the players over time.
-(b) Token allocations in rounds 1-4.
-Red dashed arrows indicate tokens used to attack while black solid arrows indicate tokens given.
-(c) Token allocations in rounds 5-6.
-To increase understanding of JHG dynamics, consider the contrived example depicted in Figure 1.
+To increase understanding of JHG dynamics, consider the example depicted in Figure 1.
 First, when players exchange (give) tokens with each other, they rise in popularity (e.g., players p2 and p3).
@@
 Finally, attacks that are stronger than the defense (strength is measured as tokens allocated multiplied by player popularity) result in inflicted damage.
-For example, p5’s attack, combined with p6’s, on p1 in rounds 5-6 is substantially stronger than p1’s defense.
+For example, p5’s and p6’s combined attack on p1 in rounds 5-6 is substantially stronger than p1’s defense.
 Thus, p1 loses popularity and p5 and p6 profit somewhat.
-While the contrived example presented above deals primarily with pairwise interactions for illustrative purposes, these dynamics can give rise to complex group behavior and emergent social norms.
+While this contrived example deals primarily with pairwise interactions for illustrative purposes, these dynamics can give rise to complex group behavior and emergent social norms.
 Frequently, alliances form and dissolve as players join together to help each other, punish those that violate norms, mitigate powerful enemies, and exploit others.
-3 Modeling human behavior in the JHG using human data
-In this section, we introduce four algorithms for learning models of human behavior in the JHG from a relatively small data set of games played by experienced human players.
-These algorithms vary based on assumptions encoded in their parameter-based representations and their methods for learning parameterizations.
+The complex strategic nature of the JHG gives rise to questions about how humans play it.
+Unlike many social dilemmas, players in the JHG must go beyond deciding whether to cooperate or defect.
+They must decide with whom they should cooperate (give tokens), when to create new relationships, when to dissolve friendships, when to defend one’s self (by keeping tokens), and who they should attack (take tokens) and when.
+In the next section, we describe methods that attempt to encode human behavior in the JHG.
+3. Modeling human behavior in the JHG
+In this section, we introduce four algorithms that learn models of human behavior in the JHG by observing the play of experienced humans.
+These algorithms vary based on assumptions encoded in their parameter-based representations and methods for learning parameterizations.
 We discuss these design choices in turn.
-3.1 Parameter-based models of behavior
-We construct algorithms based on two parameter-based models of behavior for playing the JHG: Behavior Matching (TFT) and community-aware behavior (CAB).
-Behavior Matching (TFT): In behavior matching, or tit-for-tat (TFT), a player reciprocates the actions that other players directed towards them in the previous round [4].
+3.1. Parameter-based models of behavior
+We could potentially learn models of human behavior in the JHG using a variety of approaches.
+For instance, we could take a sort of tabula rasa approach, in which we feed large amounts of human data into a neural network.
+Unfortunately, obtaining large amounts of data of people playing the JHG (or some other strategic network game) is difficult.
+Our data set currently contains only a few hundred human player games, and it is unclear how to create human-like data synthetically.
+Thus, using a data hungry approach, such as a graph neural network, appears to be somewhat impractical.
+Additionally, we would like our model to be interpretable, meaning that it can inform us about the strategies used by human players.
+Black-box methods, such as graph neural networks, typically do not provide this desired interpretability.
+As such, we consider the ability of parameter-based models to learn human behavior in the JHG.
+Such models are based on assumptions of how humans behave, with parameters that modulate behavior.
+Formally, let $m$ be the number of parameters in the model.
+Then, $\Theta=(\theta_{1},\theta_{2},\cdots,\theta_{m})$ is a parameterization of the model in which $\theta_{i}$ is the value of the $i^{\rm th}$ parameter.
+In this paper, we consider two models: (1) Behavior Matching (or tit-for-tat; TFT) axelrod1984evolution and (2) Community-Aware Behavior (CAB) skaggs2024fostering.
+Behavior Matching (TFT): In this model, a player reciprocates the actions that other players directed towards them in the previous round axelrod1984evolution.
 We consider this model due to its prevalence in prior work.
-Tit-for-tat and its variants have repeatedly been shown to be effective and robust in repeated prisoner’s dilemmas (e.g., [3, 32, 19]), a common domain for studying when and how humans and other agents cooperate with each other.
-Defining TFT in the JHG is not straightforward, as a player may receive more or less in a round than it has the ability to reciprocate.
-In such cases, an agent must determine who to exclude (when receiving too much) or what to do with extra tokens (when not receiving enough).
-We have defined a TFT algorithm for the JHG whose behavior is tuned with seven parameters.
-These parameters impact how the agent allocates tokens in the first round, what it seeks to match (token allocations or influence), and how its deals with excess and shortfalls.
-Algorithmic details are given in SM-B.
-Community-Aware Behavior (CAB): The second parameter-based model we consider is the CAB (Community Aware Behavior) algorithm [41].
-While behavior matching assumes humans primarily reciprocate the actions of others, the CAB agent is based primarily on group dynamics.
+TFT and its variants have repeatedly shown to be effective and robust in repeated prisoner’s dilemmas (e.g., axelrod1980effective; nowak1992tit; grant2013give; nowak1992evolutionary), a common domain for studying when and how humans and other agents cooperate with each other.
+Furthermore, TFT and its variants are frequently used to describe (successful) human behavior (e.g., axelrod1984evolution; grant2013give).
+As such, we consider whether human players use TFT-like strategies in the JHG, and, thus, whether a parameterized TFT algorithm can be tuned to successfully mimic human behavior.
+However, because a player’s resources are limited in the JHG, defining TFT in this domain is not straightforward, as a player may receive more or less in a round than it has the ability to reciprocate.
+In such cases, an agent must determine what to do with extra tokens (when receiving too much) or who to exclude (when not receiving enough).
+We have defined a TFT algorithm for the JHG whose behavior is tuned with seven parameters that determine how the agent allocates tokens in the first round, what it seeks to match (token allocations or influence), and how it deals with excess and shortfalls.
+Details are given in the supplementary material (SM-2).
+Community-Aware Behavior (CAB): The second parameter-based model we consider is the CAB (Community Aware Behavior) algorithm (skaggs2024fostering).
+While using behavior matching to model humans assumes that humans cooperate and attack each other due to reciprocation, using the CAB algorithm assumes that humans help and attack each other based on group dynamics moffett2019human.
+Players in the same group support (give to) each other, work together to exploit other groups, and defend each other against the attacks of outsiders.
+CAB agents follow a parameterized algorithm for how humans form, join, and modify these groups over time.
 A CAB agent determines token allocations in a round using a three-step process.
@@
 Finally, it selects token allocations designed to build up its selected group, establish itself in the group, and deal with out-of-group threats.
-CAB behavior is defined by thirty parameters which determine how it identifies existing groups, chooses a group to try to be part of, and allocates tokens to build its group.
-We use the CAB implementation provided by Skaggs et al.
-[41].
-Assuming that human behavior in the JHG can be modeled using either of these algorithms is most certainly incorrect at some level.
-However, identifying the extent to which these models can and cannot approximate human behavior can shed valuable insights into human networks.
-3.2 Learning parameterizations
-To determine how well either parameter-based model (TFT or CAB) can model human behavior in the JHG, we learn parameterizations of each model from a limited amount of human data using two different methods: Particle Swarm Optimization (PSO) [25] and Evolutionary Population Distribution Modeling (EPDM), a methodology proposed in this paper.
-Both algorithms take as input a parameter-based model and output one or more parameterizations of the model.
-Let $m$ be the number of parameters in the desired model.
-Then, $\Theta\in(\theta_{1},\theta_{2},\cdots,\theta_{m})$ is a parameterization of the model in which $\theta_{i}$ is the value of the $i^{\rm th}$ parameter.
+CAB behavior is defined by thirty parameters which determine how it performs each of these steps.
+We use the CAB implementation provided by skaggs2024fostering.
+Assuming that human behavior in the JHG can be modeled using either of these parameterized algorithms (TFT or CAB) is most certainly incorrect at some level.
+However, identifying the extent to which these models can and cannot approximate human behavior can shed valuable insights into human networks and how human behavior can be effectively simulated in them.
+3.2. Learning Parameterizations
+To determine how well TFT and CAB parameter models can model human behavior in the JHG, we learn parameterizations of each model from human data using two different methods: Particle Swarm Optimization (PSO) kennedy1995particle and Evolutionary Population Distribution Modeling (EPDM), a methodology proposed in this paper.
+Both algorithms take as input a parameter-based model $\mathcal{M}$ and output one or more parameterizations of that model.
 Particle Swarm Optimization (PSO): In PSO, a set of particles (i.e., parameterizations) are first randomly generated.
-Over time, these parameterizations are adapted to both explore the parameter space and converge toward the best parameterization.
-In our case, the best parameterization is determined by minimizing the error, over all data samples, between the token allocations made by human players and the token allocations made by the model (given the parameterization specified by the particle) in the same situations.
-In this way, PSO produces a parameterization that describes mean behavior in the data set.
-We experimented with two different error (or distance) functions: (1) mean squared error and (2) a custom built function in which the error is based on how well token allocations match across a variety of different properties (defined in SM-D).
-Because the second function worked best in all cases, we used that error function in all of our experiments.
-Evolutionary Population Distribution Modeling (EPDM): Because mean human behavior, when followed by all individuals, does not necessarily produce human-like population dynamics or realistic individual behavior, we introduce a second method, called EPDM, for modeling human behavior from data.
-EPDM attempts to model the distribution of human behavior, rather than mean behavior.
-The EPDM algorithm, outlined in Algorithm 1, receives as input a parameter-based behavior model $\mathcal{M}$ (TFT or CAB) and a set of player games $\Gamma$ (from the training data).
-Each player game $\gamma\in\Gamma$ specifies the token allocations made by a human player in each round $t$ of a game, along with the game state available to the player at the beginning of each round (specifically, ${\mathcal{P}}(t)$, ${\mathcal{I}}(t)$, and $X(t)$).
-The algorithm outputs a set of $N$ (we used 100) parameterizations of $\mathcal{M}$ that best estimate the distribution of player strategies in $\Gamma$.
-Algorithm 1 EPDM – Computes $N$ parameterizations to estimate a distribution of strategies
+Over time, these particles are adapted in a way that both explores the parameter space and converges toward the best parameterization, or the parameterization that minimizing the error, over all data samples, between the token allocations made by human players in the data and the token allocations made by the model (given the parameterization specified by the particle) in the same situations.
+PSO produces a parameterization designed to describe mean human behavior represented in the data.
+We experimented with two error (or distance) functions: (1) mean squared error and (2) a custom built function which measures how well a token allocation profile matches a target profile across a variety of properties.
+Because the second function outperformed the first in all cases, we used it in our experiments.
+This function scores the degree to which token allocation profile $\mathbf{b}$ matches the target profile $\mathbf{a}$ as follows:
+${\mathcal{S}}(\mathbf{a},\mathbf{b})={\mathcal{S}}^{+}(\mathbf{a},\mathbf{b})+{\mathcal{S}}^{-}(\mathbf{a},\mathbf{b})+{\mathcal{S}}^{\rm keep}(\mathbf{a},\mathbf{b})-{\mathcal{P}}(\mathbf{a},\mathbf{b}).$ (1)
+Here, ${\mathcal{S}}^{+}(\mathbf{a},\mathbf{b})$ scores the degree to which $\mathbf{b}$ (a) has positive allocations to the same number of players as $\mathbf{a}$, (b) has positive allocations to the same players as $\mathbf{a}$, (c) matches the total number of tokens allocated for giving as $\mathbf{a}$, and (d) gives the same number of tokens to each player as $\mathbf{a}$.
+${\mathcal{S}}^{-}(\mathbf{a},\mathbf{b})$ scores the degree to which $\mathbf{b}$ (a) has a negative allocation if $\mathbf{a}$ has a negative allocation, (b) has the same number of tokens used for taking as $\mathbf{a}$, and (c) takes tokens from the same players as $\mathbf{a}$.
+${\mathcal{S}}^{\rm keep}(\mathbf{a},\mathbf{b})$ compares the number of tokens kept in allocations $\mathbf{a}$ and $\mathbf{b}$.
+Finally, ${\mathcal{P}}(\mathbf{a},\mathbf{b})$ is a penalty term when allocation $\mathbf{b}$ gives to some player $i$ but allocation $\mathbf{a}$ takes from player $i$ or vice versa.
+Details related to ${\mathcal{S}}(\mathbf{a},\mathbf{b})$ are given in SM-4.
+Evolutionary Population Distribution Modeling (EPDM): Because mean human behavior, when followed by all individuals, may not produce human-like population dynamics or realistic individual behavior, we introduce a second method, called EPDM, for modeling the distribution of human behavior.
+The EPDM algorithm, outlined in Algorithm 1, receives as input a parameter-based behavior model $\mathcal{M}$ (we consider TFT and CAB in this paper) and a set of player games $\Gamma$ (from the training data).
+Each player game $\gamma\in\Gamma$ specifies the token allocations made by a human player in each round $t$ of a game, along with the game state available to the player at the beginning of that round.
+This game state consists of a vector ${\mathcal{P}}(t)$ specifying the current popularity of each player, a matrix ${\mathcal{I}}(t)$ specifying the influence that the players have on each other, and a matrix $X(t)$ specifying the token allocations made by all players in that round.
+The algorithm outputs a set of $N$ parameterizations of $\mathcal{M}$ that best estimate the distribution of player strategies in the player games $\Gamma$.
+We used $N=100$.
+Algorithm 1 EPDM – Computes NN parameterizations to estimate a distribution of strategies
 Input: $\mathcal{M}$ (parameter-based model), $\Gamma$ (player games), and $G$ (# generations)
@@
   (1) $\forall\pi\in\Pi(g)$, compute $\Gamma^{\pi}(g)$ (player games in which $\pi$ is a top performer)
-     (2) $\Phi(g)=$ FindCoreSet$(\Gamma,\Pi(g),\forall{\pi\in\Pi(g)},~{}\Gamma^{\pi}(g))$
+[…]
   (3) Construct $\Pi(g+1)$:
@@
 This set, denote $\Gamma^{\pi}(g)$, is determined as follows.
-Let $a^{\gamma}(t)$ denote the token allocations made in player game $\gamma\in\Gamma$ in round $t$.
-Also, let $a^{\pi}(t)$ be the allocations that parameterization $\pi$ would make in round $t$ given the history of game transactions made prior to round $t$.
+Let $\mathbf{a}^{\gamma}(t)$ denote the token allocations made in player game $\gamma\in\Gamma$ in round $t$.
+Also, let $\mathbf{a}^{\pi}(t)$ be the allocations that parameterization $\pi$ would make in round $t$ given the game history (prior to round $t$).
 Then, the average error of $\pi$ in player game $\gamma$ is
-$\delta^{\pi}_{\gamma}=\frac{1}{T_{\gamma}}\sum_{t=1}^{T_{\gamma}}dist(a^{\gamma}(t),a^{\pi}(t)),$ (1)
-where $T_{\gamma}$ is the number of rounds in player game $\gamma$ and $dist(a,b)$ is the error function specifying the distance between allocations $a$ and $b$ (see SM-D for details).
-$\pi$ is a top-performing parameterizations for player game $\gamma$ if $\delta^{\pi}_{\gamma}-\min_{\pi^{\prime}\in\Pi}\delta^{\pi^{\prime}}_{\gamma}<\varepsilon$.
-We used $\varepsilon=0.01$.
-In step 2 of Algorithm 1, a core set of parameterizations $\Phi(g)\in\Pi(g)$ is computed using the greedy process defined in the function FindCoreSet (see SM-C.2).
-FindCoreSet repeatedly adds the parameterization in $\Pi(g)$ that is a top-performer in the most player games that have not yet been represented until all player games are represented by at least one top-performing parameterization.
-Finally, in step 3, the EPDM algorithm constructs the set of parameterizations used in the subsequent generation.
-First, each parameterization from the core set $\Phi(g)$ is added to $\Pi(g+1)$.
-The remaining parameterizations then are generated using genetic evolution on the parameterizations in $\Pi(g)$.
-That is, parameterizations are evolved using selection, crossover, and mutation (see SM-C.3 for details).
-In this process, the fitness of parameterization $\pi$ is the cardinality of $\Gamma^{\pi}(g)$ (i.e., $|\Gamma^{\pi}(g)|$.
-After all generations are completed, the final distribution of parameterizations returned by the EPDM algorithm is determined by resampling $\Pi(G-1)$ based on fitness. See SM-C.4 for details.
-4 Experiment A: Evaluating population dynamics
-Our goal is to find models of human behavior in the JHG that satisfy two criteria.
-First, the population dynamics of agent collectives should mirror the population dynamics of human societies.
-Second, individual agents should display human-like behavior.
-In this section, we evaluate four models, created by combining parameter-based models (hTFT and hCAB111hCAB and hTFT denote parameterizations of CAB and TFT models, respectively, learned from human data.) with modeling algorithms (PSO and EPDM), with respect to the first criteria.
-In Section 5, we address the second criteria.
-4.1 Experiment design
-We simulated societies of hTFT-PSO, hTFT-EPDM, hCAB-PSO, and hCAB-EPDM agents in the JHG, and compare them to human societies.
-We also consider two baseline agents: (1) Random, which randomly allocates tokens with percentages of give, take, and keep that match the percentages used in the training set (Section 4.2) and (2) eCAB agents [41].
-eCABs are CAB agents with parameterizations learned via evolutionary simulations rather than from human data.
-eCABs display somewhat effective behavior in the JHG, though their strategies are distinct from humans [41].
-To evaluate these agents, we compare their population dynamics with those of human players observed in the 15 games in our test set (Section 4.2).
-Simulations for agent populations were conducted with identical settings (initial popularities, number of agents, and game length) as games found in the test set.
-For each game in the test set, four simulations of each agent population were conducted with identical numbers of agents and rounds, bringing the total number of simulations for each agent to 60.
-In setting up experiments in this way, we can directly compare human and agent societies.
-For simulations with hTFT-EPDM agents, hCAB-EPDM agents, and eCAB agents, parameterizations were randomly selected from the learned pool of strategies for each game.
-On the other hand, for hTFT-PSO and hCAB-PSO, the parameterizations to which the PSO algorithm converged from four runs of the PSO algorithm were used.
-4.2 Data sets
-Our data set, which is provided in the supplementary material, consists of 66 games played primarily by experienced human players.
-47 of these 66 games had only human players.
-The other 19 games were played by both human and bot players.
-Of the 47 all-human games, we randomly selected 15 of these games for the test set.
-The remaining 51 games were designated as training data.
-Games in the training set had 5-12 players with an average of 8.08 players (of which 6.57 were humans).
-Across these games, we have a total of 335 human player games (which comprise $\Gamma$).
-The average number of rounds in these player games was 24.96.
-Games in the test set were played by 6-11 human players (average was 7.93 players).
-Games with more than 30 rounds were truncated at 30 rounds.
-4.3 Metrics to evaluate population dynamics
-Table 1: Metrics used to compare the population dynamics of human and agent societies.
-Category Description Metrics
-Wealth & Power Wealth and influence of society members Mean popularity, Gini Index
-Economic The high-level actions (give, take, keep) % give, % take, % keep,
-Behavior taken by individuals in society Evolution Coefficient
-Mixing Connections across society, including Reciprocity, Density,
-Patterns dyadic, triadic, and group relationships Entropy, Polarization
-Summary Comparison to humans across all metrics Mahalanobis distance
-We evaluate population dynamics using the metrics summarized in Table 1.
-These metrics are grouped into four categories, each designed to evaluate a particular aspect of population dynamics.
-Metrics in the first category quantify the wealth and power wielded by society members.
-We use Mean Popularity, an approximation of Katz centrality [24], and Gini Index, which measures the distribution of wealth across society.
-Economic behavior summarizes player actions across society, including the percentages of token allocations of each type (keep, take, and give) and the average change in playar token allocations from round to round (Evolution Coefficient).
-Next, we assess societal mixing patterns, or the way that players both join together in pairs and groups, and fragments and segregates.
-These mixing patterns are quantified using two measures of reciprocity, density, entropy, and polarization.
-Finally, we summarize the dissimilarity between agent and human populations across all these metrics using Mahalanobis distance [27].
-Formal definitions of each metric, as well as motivation for why we chose these (as opposed to other) metrics, are provided in SM-E.
-4.4 Results
-Figure 2 displays measures of human and agent population dynamics. We discuss each by category.
-(a)                                     (b)                                     (c)
-Society % Give % Keep % Take
-Humans   75 19   6
-Random   72 24   4
-eCAB   95   5   0
-hTFT-PSO   97   3   0
-hTFT-EPDM   70 15 15
-hCAB-PSO 100   0   0
-hCAB-EPDM   74 19   7
-(d)                                   (e)                                       (f)
-Figure 2: Comparisons of human and agent societies. Dots (a-c), numbers (d), and lines (e-f) denote the mean, error ovals (which do not account for covariance) and ribbons show the standard error.
-Wealth and power: Figure 2a shows a comparison of agent and human populations with respect to Mean Popularity and Gini Index.
-In these two regards, hCAB-EPDM populations most closely mirror those of humans.
-Random populations tend to have both low Gini Index and low mean popularity.
-hTFT-EPDM populations tend to have slightly higher Gini Index and lower popularity, while hTFT-PSO, hCAB-PSO, and eCAB populations all have higher mean popularity and lower Gini Index.
-While high average popularity and lower Gini Index is desirable (though this does not necessarily indicate robust behavior in general), our objective is to identify models that mimic human behavior.
-With respect to mean popularity and Gini Index, the CAB model coupled with EPDM parameter modeling results in the agent model that most closely reflects human population dynamics.
-Economic behavior: In the JHG, wealth outcomes are the result of economic behavior.
-eCAB agents and algorithms that model parameters using PSO all have higher mean popularity and lower Gini Index than human populations.
-These agents use most of their tokens to give to each other.
-They keep and take little, if any (Figure 2d).
-On the other hand, humans and agents modeled with EPDM have higher variability in strategies across players and tend to keep and take more tokens.
-This produces greater inequality in the society and less overall prosperity.
-As such, agents produced using EPDM tend to make more human-like distributions of token allocations.
-The evolution coefficient measures the consistency of player token allocations over rounds.
-Figure 2e shows that none of the agent populations display similar behavior to humans in this regard.
-CAB agents are too consistent in their allocations (though hCABs are less consistent than eCABs), showing less change in token allocations from round to round than humans.
-On the other hand, TFT agents display a sort of multiple personality disorder.
-The act of reciprocating actions from the previous round produces oscillating changes in token allocations that are clearly not human-like.
-However, the differences in token allocations of TFT agents over two rounds are quite small.
-Mixing patterns: We measure two forms of reciprocity: Immediate Reciprocity (reciprocation from one round to the next) and Overall Reciprocity (reciprocation over a complete game).
-In the JHG, humans have moderate levels of both forms of reciprocity (Figure 2b).
-EPDM agents exhibit similar levels of reciprocity, with hCAB-EPDM’s rates being only slightly higher than people’s.
-On the other hand, eCABs and PSO reciprocation rates are far higher than those of humans (in the data set).
-The density and entropy measures, shown in Figure 2c, provide information about the number (density) and strength (entropy–are all trading partners valued similarly?) of trading partners.
-In these regards, hCAB-EPDM agents once again tend to be very similar to those of human players.
-hCAB-PSO agents have similar density, but higher entropy.
-On the other hand, agents using TFT tend to be less similar to humans in both regards.
-Thus, hCAB agents more effectively mirror these human behaviors better than TFT agents, with hCAB-EPDM being the best.
-Finally, we consider polarization (Figure 2f), which describes the degree to which players segregate into groups.
-Since the CAB algorithm is based on group formation, while TFT is not), it is not unexpected to observe that all CAB agents display higher polarization than TFT agents.
-These higher levels are similar to the polarization measures we observe in human games.
-This again shows that the CAB model can be tuned to better model human dynamics than the TFT (matching) model.
-Metric Random eCAB hTFT-PSO hTFT-EPDM hCAB-PSO hCAB-EPDM
-Mahalanobis Dist. 18.822   7.089 11.565   7.656   9.216   3.393
-Chi-Square Test $p<0.001$ $p<0.001$ $p<0.001$ $p<0.001$ $p<0.001$ $\mathbf{p=0.486}$
-Table 2: Mahalanobis distance (lower is better) between human and agent populations over all metrics. $p>0.05$ indicates that no difference is detected between human and agent populations.
-Summary metric: Table 2 gives the Mahalanobis distance between the various agent populations and our human population across the eleven individual metrics.
-Of the six agent populations, hCAB-EPDM populations most closely mirror human societies from our test set.
-Indeed, a Chi-Square test detects no statistical difference between hCAB-EPDM and human populations ($p=0.486$), thus indicating a relatively close fit across these measures.
-All other agent populations were statistically distinct from human populations ($p<0.001$ in all cases).
-These results show that the CAB algorithm paired with EPDM parameterization modeling produces agent populations whose dynamics are surprisingly close (with some variation–notably the evolution coefficient) to those of human populations in the JHG (first criteria).
-In the next section, we consider whether individual hCAB-EPDM agents plausibly exhibit human behavior (second criteria).
-5 Experiment B: Interacting with human players
-In this section, we seek to establish the degree to which individual hCABs behaviors seem human-like to people. To do this, we conducted a user study pairing human players with hCABs in the JHG.
-5.1 User study design
-We recruited eight people, each of which had prior experience playing the JHG, to participate in the study.
-The study proceeded in a series of four periods.
-In each round, the participants were randomly divided in half.
-Each group of four participants was paired with four (randomly selected) hCABs to form two games played by eight players each.
-Thus, eight games were played across the four periods.
-Games lasted for 21, 20, 17, and 15 rounds in Periods 1-4, respectively (players were not told game lengths).
-For consistency, only the first 15 rounds of each game were used in our data analysis.
-Participants were asked to play so as to become as popular as possible by the end of each game, while also identifying whether associates were humans or bots.
-Participants were told in advance that each game was composed of four people and four bots.
-After each game, participants completed a survey asking them to identify each associate as a human or a bot.
-We also compare humans and hCABs with respect to their popularity and the percentage of tokens used to give, keep, and take.
-5.2 Results
-Society % Give % Keep % Take
-Humans   60 25 15
-hCABs   75 20   5
-(b) Action percentages
-                                                                      Guesser % Correct Designation Humans 56.7 Chance 57.1
-                  (a) Popularity over time                           (c) Detecting humans vs. hCABs
-Figure 3: User study results comparing humans and hCABs.
-(a) Mean popularity over time.
-Error ribbons show the standard error.
-(b) Percent of tokens used to give, keep, and take by each player type.
-(c) Percentage of players correctly identified by human participants verses random chance.
-Over the eight games, hCABs had slightly higher average popularity than humans in rounds 5-15 (Figure 3a).
-This was potentially caused by some human participants being quite aggressive, attacking substantially more than hCABs (Figure 3b).
-Attacks sometimes lead to retaliation and prolonged fighting, resulting in lost popularity.
-While the distribution of human actions was different than humans in our data sets (see previous section), average hCABs behavior remained about the same.
-Despite these differences between these particular human participants and hCABs, participants were unable to distinguish each other from hCABs on average.
-Figure 3c shows that participants correctly identified whether their associates were human or bot 56.7% of the time.
-Given that participants knew that four of their associates were bots and three were humans, random chance would have produced a correct detection rate of 57.1%.
-Therefore, participants were no better than random in identifying whether there associates were humans or hCABs.
-This gives credence to the argument that the behaviors of the randomly selected hCABs seem human-like.
-Viewing hCAB behavior in individual games provides further evidence of this claim.
-For example, consider Figure 4, which depicts one of the eight games from the study (similar graphs for the other games are provided in SM-G).
-In Round 9, two human players (p1 and p2) attacked each other.
-Interestingly, three of the four hCABs and the two other humans all joined with p7 in attacking p1 in Round 10 (which greatly weakened p7 for the remainder of the game).
-That the hCABs all chose to take the same side in the dispute between p7 and p1 as the other human players (siding with the better connected individual) illustrates the ability of hCABs to reason about groups and power in a way that led them to mirror human behavior in this situation.
-Similar observations can be made about how hCABs joined together in subgroups with humans and other hCABs in these games.
-Figure 4: Illustrative game showing the popularity of players over time (top) and network structure in select rounds (bottom).
-In the graphs, arrows indicate token transactions in the current round (green = give; red = take), while nodes more connected historically tend to be closer to each other.
-6 Conclusions and delimitations
-In this paper, we compared and contrasted several different algorithms in modeling human behavior in a strategic network game called the Junior High Game (JHG) from a relatively small data set.
-We considered approaches that differ with respect to the underlying assumptions of human behavior they use (behavior matching verses community-aware behavior) and the statistical moments they model (mean verses distribution).
-Our results indicate that a parameter-based community-aware model (CAB [41]) coupled with evolutionary population distribution modeling (EPDM) can produce agents that closely reflect human population dynamics in the JHG.
-Furthermore, results of a user study indicate that the individual behaviors of these agents plausibly mimic human decision-making.
-The CAB model consists of a variety of hand-coded reasoning functions modulated by thirty parameters.
-Such a modeling approach has advantages and disadvantages compared to a more tabula rasa approach (such as a graph neural network).
-One advantage is that an effective model of human behavior can be derived using only a small data set.
-This is valuable, since obtaining large amounts of quality data in strategic network games can be challenging.
-A second advantage is that the trained strategies are interpretable, thus potentially shedding insight on human behavior.
-On the other hand, a tabula rasa approach could potentially learn other kinds of strategies, and is therefore a potentially promising area for future work.
-hCAB-EPDM provides a baseline for such future methods.
-While our results indicate that hCAB-EPDM agents are, perhaps surprisingly, effective models of human behavior in the JHG, we caution against overly broad extrapolations.
-For example, the societies modeled in this paper were small (6-11 players).
-Larger societies have different characteristics.
-Additionally, the hCAB-EPDM agents used in this paper were trained on a small data set.
-Given the complex state and strategy space of the JHG, other human players could potentially use different strategies not found in this data set.
-Though the CAB algorithm can be configured (via its parameters) to model many different kinds of behavior, it is possible that other communities of human players could display vastly different norms of acceptable behavior that may not be effectively modeled by parameterizations of the CAB algorithm.
-Future work could consider whether behavior in larger, more varied, data sets could be modeled effectively by CAB agents.
-Finally, the value of the results described in this paper somewhat hinges on the degree to which the strategic network game we selected (the JHG) models human networks.
-While the JHG models important properties of such networks, the world is certainly more complex than any game can model.
-Future work should consider refinements and enhancements to study other complexities of human networks.
-These and other future works can potentially help us continue to better understand and model human networks.
-Appendix A Overview of the Supplementary Material
-The supplementary material consists of two parts:
-  1. 1.
-    Code and data sets used on the study. These are supplied in a zip file.
-  2. 2.
-    The additional explanations and results provided below that support the explanations and results contained in the main paper.
-The additional explanations that follow cover six topics:
-  1. 1.
-    Section B: Description of the parameter-based TFT algorithm.
-  2. 2.
-    Section C: Additional information about the EPDM algorithm.
-  3. 3.
-    Section D: Additional information about the error function used by the EPDM and PSO algorithms.
-  4. 4.
-    Section E: A formal description of the metrics used to evaluate the population dynamics of human and agent populations.
-  5. 5.
-    Section F: Additional notes about the first experiment reported in the main paper.
-  6. 6.
-    Section G: Additional notes and results related to the second experiment (user study) reported in the main paper.
-Appendix B A parameterized-based tit-for-tat agent for the JHG
-In this section, we defined the parameter-based matching algorithm (tit-for-tat; TFT) discuss in this paper to model human behavior in the JHG.
-Reciprocating the behavior of associates in the JHG is not as straightforward as it is in repeated games.
-A TFT algorithm for the JHG must specify three things.
-First, the algorithm must specify how the player will allocate tokens in the first round.
-Second, the algorithm must determine what the player will reciprocate.
-A player could reciprocate based on either the number of tokens given (or taken) or the amount of influence (number of tokens multiplied by popularity) given (or taken).
-Third, the algorithm must determine what to do when agents receive more or less tokens from other players than they are able to reciprocate.
-The seven parameters specified in Table 3 dictate the token allocations made by the agent.
-In the hTFT agents used in this paper (hTFT-PSO and hTFT-EPDM), these parameters are tuned to match behavior specified in the training set as closely as possible using the PSO and EPDM modeling methods, respectively.
-We define how TFT agents act given these parameters in this section. The algorithm is implemented in the file TFTAgent.h in the supplied code.
-Table 3: The set of parameters $\Theta$ used in the TFT parameter-based algorithm. Each parameter can take on an integer value in the range $[0,100]$.
-Symbol Usage
-$\theta_{\rm initKeep}$ Specifies the percentage of its tokens the agent keeps in the first round
-$\theta_{\rm initAllocSize}$ Specifies the average number of tokens allocated to selected associates in the
-first round
-$\theta_{\rm initPercNeg}$ Specifies the probability that an initial token allocation is positive (give) or
-negative (attack)
-$\theta_{\rm matchType}$ Determines whether the agent matches tokens ($<50$) or influence ($\geq 50$)
-$\theta_{\rm notEnough}$ Determines, in the case of not being able to reciprocate everything, whether the
-agent scales back all token allocations proportionally ($<50$) or prioritizes higher
-impact relationships ($\geq 50$)
-$\theta_{\rm tooMany}$ Determines what to do in case of having excess tokens. When $<50$, the agent
-scales up all token allocations proportionally. Otherwise, it creates new random
-allocations.
-$\theta_{\rm keepExtra}$ Specifies the percentage of extra tokens (after reciprocating everything) that the
-agent keeps
-B.1 First-round token allocations
-The first three parameters in Table 3 ($\theta_{\rm initKeep}$, $\theta_{\rm initAllocSize}$, and $\theta_{\rm initPercNeg}$) impact how the agent allocates tokens in the first round.
-The following rules define the agent’s behavior in this round:
-  • $\theta_{\rm initKeep}$ specifies the percentage of tokens the agent keeps in the first round.
-That is, the agent keeps $\left(\frac{N\cdot\theta_{\rm initKeep}}{100}\right)$ tokens, rounded to the nearest integer, in the first round.
-Here, $N$ is the number of tokens it has to allocate.
-We used $N=2|I|$ (two times the number of players) in our games.
-  • The remaining tokens are allocated by randomly selecting players in the game (who will be the recipient of an allocation) and then allocating a subset of tokens to them.
-The number of tokens allocated to the selected agent is $z=1+{\rm round}\left((N-1)*\frac{R}{100}\right)$, where ${\rm round}(\cdot)$ rounds the value to the nearest integer, $R=\max(0,\min(100,\theta_{\rm initAllocSize}+r_{20}-10$)), and $r_{20}$ is an integer randomly selected in the interval $[0,20]$.
-If $z$ is greater than the number of tokens the agent has remaining to allocate, it is truncated to the remaining number of tokens.
-  • The allocations made in the previous bullet point can be positive (give) or negative (attack).
-$\theta_{\rm initPercNeg}$ specifies the probability that the agent makes each allocation positive or negative.
-B.2 Determining what to reciprocate
-In the JHG, there are two ways one can think about matching behavior.
-First, the agent can just reciprocate the number of tokens (positive or negative) that a player allocated to them in the previous round.
-However, because the tokens of more popular players are more influential than the token allocations of less popularity people, an agent could also choose to match influence instead of number of tokens.
-The influence exchanged in a token allocation is give by the popularity of the allocator $P_{i}$ multiplied by the number of tokens allocated.
-$\theta_{\rm matchType}$ specifies whether the TFT agent matches tokens or influence.
-If $\theta_{\rm matchType}<50$, then the agent matches tokens.
-Otherwise, it matches influence.
-B.3 Dealing with excess and shortfall
-Regardless of what the agent chooses to reciprocate (tokens or influence), it can receive more or less in a round than it has the ability to reciprocate. We deal with each case individually.
-First, in the case of receiving more than it can reciprocate, the agent uses one of two methods.
-When $\theta_{\rm notEnough}<50$ the agent simply computes the number of tokens needed to match all transactions from each agent, and then scales back all allocations proportionally.
-When $\theta_{\rm notEnough}\geq 50$, the agent first reciprocates fully with agents with whom it has historically had the most interactions, determined by $|{\mathcal{I}}(i,j)|+|{\mathcal{I}}(j,i)|$, where $|{\mathcal{I}}(i,j)|$ is the influence of player $i$ on player $j$ in the JHG influence matrix [41].
-It continues to fully reciprocate to those agents until it has no remaining tokens.
-Thus, it will not tend to reciprocate the actions of associates with whom the overall strength of its interactions have been small in the past.
-Second, in the case of having more tokens than are needed to reciprocate all of last round’s transactions, the agent first ensures that it reciprocates all allocations made to it in the last round by other players.
-It then keeps a percentage of the remaining tokens specified by $\theta_{\rm keepExtra}$.
-Finally, with the $100-\theta_{\rm keepExtra}$ percent of the remaining tokens, it simply scales up all token allocations proportionally when $\theta_{\rm tooMany}<50$, or, otherwise, randomly allocates tokens (positively or negatively based on $\theta_{\rm initPercNeg}$) to randomly selected associates.
-Appendix C EPDM
-C.1 Implementation
-The implementation of the EPDM algorithm we used is given in the file EPDM.cpp of the supplied code. Note that we used a gene pool of size $N=100$ parameterizations in each generation.
-C.2 FindCoreSet function
-Due to space limitations, the FindCoreSet function was not provided in the main paper.
-This algorithm is given in Algorithm 2 of this supplementary material.
-Variables are defined as in the main paper.
+$\delta^{\pi}_{\gamma}=\frac{1}{T_{\gamma}}\sum_{t=1}^{T_{\gamma}}dist(\mathbf{a}^{\gamma}(t),\mathbf{a}^{\pi}(t)),$ (2)
+where $T_{\gamma}$ is the number of rounds in player game $\gamma$ and $dist(\mathbf{a},\mathbf{b})$ is the error function specifying the distance between allocations $\mathbf{a}$ and $\mathbf{b}$ (we use ${\mathcal{S}}(\mathbf{a},\mathbf{b})$; Eq.
+1).
+$\pi$ is a top-performing parameterization for player game $\gamma$ if $\delta^{\pi}_{\gamma}-\min_{\pi^{\prime}\in\Pi}\delta^{\pi^{\prime}}_{\gamma}<\varepsilon$.
+We use $\varepsilon=0.01$.
+In step 2 of Algorithm 1, a core set of parameterizations $\Phi(g)\subseteq\Pi(g)$ is computed using the greedy process defined in the function FindCoreSet (see Algorithm 2).
+FindCoreSet repeatedly adds the parameterization in $\Pi(g)$ that is a top-performer in the most player-game sets that have not yet been represented until all player games are represented by at least one top-performing parameterization.
 Algorithm 2 FindCoreSet – Selects a set of parameterizations from the current set of parameterizations $\Pi(g)$ to describe the set of observed human strategies.
@@
 return $\Phi$
-C.3 Details of evolutionary algorithm
-Step 3 of the EPDM algorithm (see Algorithm 1 in the main paper) specifies that new parameterizations are added to $\Pi(g+1)$ (the parameterization/gene pool for the next generation) using genetic evolution.
-As in standard of genetic evolution, this algorithm uses selection, crossover, and mutation.
-This was done as follows:
-  1. 1.
-    Two parameterizations are randomly (based on fitness) selected from $\Pi(g)$.
-The probability a parameterization $\pi$ is selected is given by $\frac{|\Gamma^{\pi}(g)|}{\sum_{\pi^{\prime}}|\Gamma^{\pi^{\prime}}(g)|}$, where $|\Gamma^{\pi}(g)|$ is the cardinality of the set $\Gamma^{\pi}(g)$ (the number of player games for which $\pi$ is a top performer).
-  2. 2.
-    A new parameterization is formed from the two parameterizations selected in the prior step.
+Finally, in step 3, the EPDM algorithm constructs the set of parameterizations used in the subsequent generation.
+First, each parameterization from the core set $\Phi(g)$ is added to $\Pi(g+1)$.
+The remaining parameterizations are then generated using genetic evolution on the parameterizations in $\Pi(g)$ using selection, crossover, and mutation.
+First, two parameterizations are randomly (based on fitness, defined as the number of player games for which the parameterization is a top performer) selected from $\Pi(g)$.
+That is, the probability parameterization $\pi$ is selected is $\frac{|\Gamma^{\pi}(g)|}{\sum_{\pi^{\prime}}|\Gamma^{\pi^{\prime}}(g)|}$, where $|\Gamma^{\pi}(g)|$ is the cardinality of the set $\Gamma^{\pi}(g)$.
+Second, a new parameterization is formed from the two parameterizations selected in the prior step.
 In this parameterization, the $i^{\rm th}$ value is determined in one of three ways.
@@
 Otherwise, the value is chosen to be the $i^{\rm th}$ value of one of the two selected parameterizations (chosen randomly).
-With probability 0.12, this value is then randomly shifted up or down from $[0,5]$ values (truncated to be between 0 and 100, inclusive).
-C.4 Resampling
-The last step of the EPDM algorithm (Algorithm 1 from the main paper) consists of resampling the parameterizations evaluated in the last generation $\Pi(G-1)$ based on fitness.
+With probability 0.12, this value is then randomly shifted up or down 0 to 5 values (truncated to be between 0 and 100, inclusive).
+After all generations are completed, the last step of the EPDM algorithm (Algorithm 1) consists of resampling the parameterizations evaluated in the last generation $\Pi(G-1)$ based on fitness.
 This is done by randomly selecting $N$ parameterizations from the list $\Pi(G-1)$.
@@
 Note that a parameterization can be chosen more than once.
-Appendix D Error Functions
-Both the PSO and EPDM algorithms for learning parameterizations from data require an error or distance function (e.g., Eq.
-1 in the main paper).
-This function takes as input two token allocation profiles, and returns a value specifying the difference between them.
-We experimented with several such functions, including mean-squared error (MSE) and a custom built function, called property scoring (PS).
-In PS, the error is based on how well the two token allocation profiles match across a variety of different properties.
-More specifically, the PS error function scores pairs of token allocation profiles using a set of properties.
-The distance between the actual allocations and the proposed allocations is decreased for keeping the right number of tokens, giving the right number of tokens in total, giving tokens to the right players, giving the right number of tokens to the right players, taking at the right times, taking the right number of tokens, and taking from the right player.
-The distance function is increased for proposing to give to a player from which tokens were actually taken or taking tokens from a player for which tokens were actually given.
-Exact specifications of the property scoring distance function are given in the file EPDM.cpp (see the function scoreProposedAllocation) of the supplied code.
-Because the PS error function always performed better than MSE, we used this scoring function in the training of all agent models.
-Appendix E Metrics
-In this section, we define how we computed each of the metrics used in the paper to evaluate population dynamics.
-These metrics are derived from the known quantities of the JHG, including the transaction matrix $X$, the popularity vector ${\mathcal{P}}$, and the influence matrix $\mathcal{I}$.
-Additionally, to capture longer-term (non-discounted) interactions between players, we use the impact matrix.
-We define the impact matrix prior to discuss each of the metrics used in the paper.
-We define the economic actions taken by a player at time $\tau$ as a matrix $X(\tau)$.
-Some metrics, discussed below, use both positive and negative connections while others only one or the other.
-To distinguish between each case while trying to simplify explanations we distinguish between $X(\tau)$, $X^{+}(\tau)$, and $X^{-}(\tau)$ as all connections, positive connections, and negative connections.
-We apply this to the different derived matrices as well, such as the influence matrix or impact matrix presented below.
-E.1 The impact matrix
-To examine the evolution of friendships and interaction patterns throughout the Junior High Game, we construct the impact graph, a network representation that aggregates player interactions over some duration of the game.
-This graph encodes social mixing patterns by weighting interactions based on their frequency, the relative popularity of the involved players at the time of interaction, and the corresponding economic action coefficient.
-To calculate the impact matrix, we first calculate $J^{\prime}_{ij}(\tau)$ representing the interaction strength between players $i$ and $j$ at time step $\tau$. It is defined as:
-$J_{ij}(\tau)=\begin{cases}\mathcal{P}(\tau)X_{ij}(\tau)c^{keep},&\text{if }i=j\\ \mathcal{P}(\tau)X_{ij}(\tau)c^{give},&\text{else if }x_{ij}\geq 0\\ \mathcal{P}(\tau)X_{ij}(\tau)c^{take},&\text{else if }x_{ij}<0\end{cases}$ (2)
-where $\mathcal{P}_{i}(\tau)$ represents the popularity of player $i$ at time $\tau$ adjusting the weight of interactions based on social influence.
-$X_{ij}(\tau)$ is the economic action between players $i$ and $j$ with $c^{give}$, $c^{keep}$, and $c^{take}$ representing the economic action coefficients that assign weights based on the type of interaction (giving, keeping, or taking) respectively.
-The aggregated impact matrix over a time interval $[q,r]$ is given by:
-$J^{\prime}(q,r)=\sum_{\tau=q}^{r}J(\tau)$ (3)
-which sums the impact matrices over all time steps within the interval. We then normalize the interaction weights so that we can compare relative influences across different context as follows:
-$J^{\prime}(q,r)=\frac{J(q,r)}{\|J(q,r)\|_{1}}$ (4)
-where $\|J^{\prime}(q,r)\|_{1}$ is the L1 norm of the matrix, ensuring that the sum of the magnitude of all entries in $J(q,r)$ is constrained within a unit scale.
-Additionally, we often look at the impact of just one round. That is, if $q=r$, we use the notation $J^{\prime}(r)$ to mean $J^{\prime}(q,r)$ to simplify understanding of our equations.
-E.2 Metrics for wealth and power
-Because the JHG uses popularity (an approximation of Katz centrality [24]) as the primary commodity, we utilize the mean popularity, $\mu_{\mathcal{P}}$, as a metric to quantify the productivity of a society.
-To apply this measure, we first constructed a popularity vector $\mathcal{P}=[\mathcal{P}_{0},\mathcal{P}_{1},\ldots,\mathcal{P}_{n-1}]$ representing the popularity of each agent $i$ at a particular time $\tau$.
-We then define average popularity as follows:
-$\mu_{\mathcal{P}}(\tau)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{P}_{i}(\tau)$ (5)
-We additionally employ the Gini Index, $G(\tau)$, as a quantitative metric to assess inequality of popularity across a society.
-The Gini Index, originally developed as a measure of income inequality [16], ranges from 0 to 1, where 0 indicates perfect equality (i.e., all individuals possess equal popularity) and 1 denotes maximal inequality (i.e., one individual possesses all of the popularity while others have none).
-$G(\tau)$ was then computed using the following standard formula:
-$G(\tau)=\frac{1}{2n^{2}\mu_{\mathcal{P}}(\tau)}\sum_{i,j=0}^{n-1}|\mathcal{P}_{i}(\tau)-\mathcal{P}_{j}(\tau)|$ (6)
-where $n$ is the number of agents and $\mu_{\mathcal{P}}$ is the mean of the values in vector $\mathcal{P}$.
-In our analysis, we used the Gini Index (as opposed to the variance of the population) because it is a preferred measure used in measuring inequality in societies in economics [38, 2] and sociology [28].
-We used average popularity and the Gini Index as an interpretable and sensitive measures for assessing the mean and spread of popularity in a society.
-E.3 Metrics for quantifying of economic behavior
-To help understand how individuals in a society interact with each other, we analyze their economic actions using the percent of give, keep and take allocations over societies of each type of agent.
-Precise definitions of these metrics are as follows:
-[…]
-[…]
-[…]
-where $X_{ij}(\tau)$ represents the token allocations of player $i$ toward player $j$ at time $\tau$.
-The diagonal of the transaction matrix $X(\tau)$ specifies agent keeping, while positive non-diagonal values represent giving and negative non-diagonal values represent taking.
-$T$ is the number of rounds in the game, $|I|$ is the cardinality of the set $I$ (specifying the number of players in the game), and $\delta$ is the Kronecker delta function.
-E.3.1 The evolution coefficient
-To capture the evolution of economic behaviors in the JHG over time, we use the evolution coefficient.
-The evolution coefficient quantifies structural changes in the impact matrix over time.
-Given that the impact matrix $J^{\prime}(q,r)$ represents normalized interaction patterns over a specified window.
-This coefficient, denoted $C_{d}$, is computed by evaluating differences in consecutive impact matrices, shifted by a delay $d$:
-$C_{d}=\sum_{\tau=0}^{r-p-1}\sum_{ij}\left|J^{\prime}_{ij}(\tau)-J^{\prime}_{ij}(\tau+d)\right|$ (10)
-In words, the evolution coefficient quantifies the cumulative absolute differences between the impact matrices over a specified number of rounds $d$.
-A higher value of $C_{d}$ indicates more substantial shifts in the interaction structure, suggesting dynamic social mixing or evolving economic strategies among players.
-Conversely, lower values suggest stability in social interactions and player behaviors.
-We note that, in the computation of the Mahalanobis distance, the evolution coefficient is measured as the average change in token allocations across one and two rounds of play.
-E.4 Mixing patterns: Measuring reciprocation
-Reciprocity is an important component of cooperative dynamics within agent societies [31].
-High reciprocity suggests that players consistently return favors and maintain balanced exchanges, whereas low reciprocity indicates asymmetric relationships, where some players give or take more than they receive.
-We define the pairwise global reciprocity between agents $i$ and $j$ over the course of an interaction as:
-$R^{global}_{ij}=\begin{cases}0&\text{if }i=j;J_{ij}(r_{0},r_{f})\leq 0;J_{ji}(r_{0},r_{f})\leq 0\\ J^{\prime}_{ji}(r_{0},r_{f})\frac{J_{ij}(r_{0},r_{f})}{J_{ji}(r_{0},r_{f})}&\text{else if }J_{ij}(r_{0},r_{f})<J_{ji}(r_{0},r_{f})\\ J^{\prime}_{ji}(r_{0},r_{f})&\text{else if }J_{ij}(r_{0},r_{f})\geq J_{ji}(r_{0},r_{f})\\ \end{cases}$ (11)
-Here, $J_{ij}(r_{0},r_{f})$ denotes the cumulative reciprocated impact from agent $j$ to agent $i$ over the interval $[r_{0},r_{f}]$, where $r_{0}$ if the first round of the game and $r_{f}$ is the final round of the game.
-The metric $R^{global}_{ij}$ thus captures the extent to which agent $j$ reciprocates the contributions made by agent $i$, normalized by the directionally lesser interaction.
-We use the overall global reciprocity coefficient to define overall reciprocity. It is summarizes the aggregate reciprocity across all agent pairs in the population as follows:
-$R^{global}_{coeff}=\frac{\sum_{ij}{R^{global}_{ij}}}{n(n-1)}$ (12)
-where $n=|I|$ denotes the number of players in the game.
-This coefficient provides a population-level estimate of the prevalence of mutual exchanges, weighted by the magnitude of impact transferred between agents over the duration of the game.
-In addition to this cumulative measure, we also define immediate reciprocity to capture short-term reciprocation dynamics. The pairwise local reciprocity at round $r_{m}$ is defined as:
-$R^{local}_{ij}=\begin{cases}0&\text{if }i=j;J_{ij}(r_{m})\leq 0;J_{ji}(r_{m})\leq 0\\ J^{\prime}_{ji}(r_{m-1})\frac{J_{ij}(r_{m})}{J_{ji}(r_{m-1})}&\text{else if }J_{ij}(r_{m})<J_{ji}(r_{m+1})\\ J^{\prime}_{ji}(r_{m-1})&\text{else if }J_{ij}(r_{m})\geq J_{ji}(r_{m-1})\\ \end{cases}$ (13)
-The corresponding local reciprocity coefficient is:
-$R^{local}_{coeff}=\frac{\sum_{ij}{R^{local}_{ij}}}{n(n-1)}$ (14)
-This formulation evaluates how frequently short-term exchanges are reciprocated from one round to the next, capturing the temporal responsiveness of agents to recent interactions.
-E.5 Mixing patterns: Measuring connectivity using density and entropy
-Graph density provides a measure of how interconnected a network is at a given time, relative to the maximum possible number of directed edges (excluding self-loops).
-Just to note, when average across the whole network average in/out degree are proportional to density and therefore we exclude the redundant comparison.
-Density is defined as:
-$D(\tau)=\frac{1}{n(n-1)}\sum_{i,j=0;i\neq j}^{n-1}A^{+}_{ij}(\tau)$ (15)
-Here, $A^{+}_{ij}(\tau)$ represents an adjacency matrix which is $1$ if a positive economic action from node $i$ to node $j$ took place in round $\tau$.
-Otherwise it is 00.
-The denominator $n(n-1)$ reflects the maximum number of actions in a loop-free graph.
-While density describes the average number of associates a player is interacting with, entropy quantifies the distribution of a player’s actions among agents.
-We define a normalized entropy score that captures how evenly distributed its outgoing edge weights are.
-For node $i$, let the set of positive outgoing edge weights be $\{X^{+}_{ij}\}_{j=1}^{n}$.
-Define the normalized edge weight distribution as:
-$w_{ij}=\frac{X_{ij}^{+}}{\sum_{k=1}^{n}X_{ik}^{+}}$ (16)
-The entropy for node $i$ is then given by:
-$H_{i}=-\sum^{n-1}_{j=0;X^{+}_{ij}>0}w_{ij}\log w_{ij}$ (17)
-The maximum possible entropy $H_{i}^{\text{max}}$ occurs when all outgoing edges from node $i$ are equally weighted, i.e.,
-$H_{i}^{\text{max}}=\log k_{i},$ (18)
-where $k_{i}$ is the out-degree of node $i$ (number of positive-weight outgoing edges). Then entropy is then computed as:
-$H_{i}^{\text{norm}}=\begin{cases}0&\text{if }\sum_{j}X^{+}_{ij}=0\text{ or }k_{i}=1\\ \frac{H_{i}}{H_{i}^{\text{max}}}&\text{otherwise}\end{cases}$ (19)
-Finally, the average entropy across all nodes provides a summary statistic for the overall spread of edge weight distributions in the network:
-$H^{\text{avg}}=\frac{1}{n}\sum_{i=1}^{n}H_{i}^{\text{norm}}$ (20)
-E.6 Measuring group formation: Polarization
-Finally, we desire to quantify how players cluster and separate.
-Known metrics for (partially) quantifying such behaviors include the clustering coefficient [47, 29] and modularity [30].
-However, many metrics either (1) are not designed for signed, weighted, and time-varying graphs, (2) do not provide comparisons of grouping behavior from graph to graph (e.g., modularity measures how well a community partition describes a graph, not how divided or grouped a graph is), or (3) conflate multiple concepts (e.g., high density tends to produce a high clustering coefficient and thus may not indicate how members of society actually group and segregate).
-Thus, in this paper, we use a customized metric called polarization to quantify how individuals separate into different groups and sub-groups.
-Related somewhat to the idea of modularity, polarization measures how much individuals would consistently prefer individuals within a designated group as opposed to individuals outside of that group.
-Polarization measures the maximum preference separation between two distinct (and sufficiently important) sub-groups in the society.
-Preference separation is a method for measuring how separated two distinct sub-groups within society are in a weighted and signed graph.
-Specifically, it measures how much individuals in a sub-group would prefer (based on immediate connections) any one member of their own sub-group to any one member of the other sub-group.
-Preference separation between two groups is highest when (1) nodes with each sub-group are well connected, (2) connections between the two sub-groups are weaker than connections within the sub-groups, and (3) the sub-groups cover a larger proportion of the entire society.
-Before giving details for how preference separation and polarization are computed, we first define several quantities. Let
-$m_{i}=\sum_{j}|\mathcal{I}_{ji}|$ (21)
-be the magnitude of influence (both positive and negative) flowing to individual $i$ from all other individuals in society. Furthermore, let
-$w_{i}=\frac{m_{i}}{\sum_{j\in I}m_{j}},$ (22)
-where $I$ is the set of all individuals in the society, be the total proportion of influence that is flowing to individual $i$.
-Next we define a set of dueling sub-groups in society by comparing how society would partition itself if all individuals had to choose between individuals $i$ and $j$ (where $i\neq j$. Let
-$G_{ij}=\{i\}\cup\{k\in I:\mathcal{I}_{ik}>\mathcal{I}_{jk}+\delta\},$ (23)
-where $\delta=0.04m_{k}$, denote the group of individuals in society that would prefer $i$ over $j$.
-Similarly, $G_{ji}$ denotes the group of individuals that prefer $j$ over $i$.
-Finally, let $G_{ij}^{\sim}=I-(G_{ij}\cup G_{ji})$ denote the set of individuals in society that are indifferent between individuals $i$ and $j$.
-Finally, let $W_{ij}=\sum_{k\in G_{ij}}w_{i}$ denote the collective weight of group $G_{ij}$.
-Preference separation between sub-groups $G_{ij}$ and $G_{ji}$, denoted $\mathcal{S}(G_{ij},G_{ji})$, measures the proportion of individual in $G_{ij}$ and $G_{ji}$ that would prefer to stay in the same group if the choice were between other individuals from the respective groups (other than just $i$ and $j$).
-Formally, it is computed as follows:
-$\mathcal{S}(G_{ij},G_{ji})=\left(\frac{\sum_{k\in G_{ij}}\sum_{\ell\in G_{ji}}|G_{ij}\cap G_{k\ell}|+|G_{ji}\cap G_{\ell k}|+0.5|G_{k\ell}^{\sim}|}{|G_{ij}||G_{ji}||I|}\right)\left(\frac{{\rm min}(W_{ij},W_{ji},0.2}{0.2}\right),$ (24)
-where ||$\cdot$|| ⋅ | denotes the cardinality of a set.
-Note that the last item in the numerator of left term gives half credit when individuals are neutral between the two groups for a particular pairing.
-The right component of the equation scales down preference separation if one of the groups is too small (determined as having a weight of less than 0.2).
-Note that measures of separation tend to be between 0.5 and 1.
-As such, we base polarization on a normalized version of preferences separation, given by $\mathcal{S}^{\prime}(G_{ij},G_{ji})=\frac{\mathcal{S}(G_{ij},G_{ji})-0.5}{0.5}$.
-Once preference separation is computed for all pairs of sub-groups, polarization is determined to be the max preference separation for pairs of subgroups.
-That is, polarization (denoted $\Gamma$) is given by $\Gamma=\max_{i,j}\mathcal{S}^{\prime}(G_{ij},G_{ji})$.
-The constraint to only consider subgroups of a particular size ensures that the preference separation between the groups could have a meaningful impact on society.
-All of the above description omit the time $t$ – polarization is computed for each time period.
-The actual metric we report is the average polarization over all rounds considered.
-We note that this metric is not computationally efficient in this formulation, but it can be computed and utilized effectively for small societies such as those we consider in this paper.
-E.7 Summary metric: Mahalanobis distance
-Each of the above metrics represent important aspects of population dynamics.
-To further compare and contrast human and agent populations, we aggregate these metrics into a multivariate Gaussian.
-That is, we quantify the dissimilarity between populations with respect to these metrics using the Mahalanobis distance [27].
-Unlike the Euclidean distance, which assumes independence across dimensions, the Mahalanobis distance incorporates the covariance structure of the data, making it well-suited for comparing complex, interdependent societal metrics.
-Given two societal types, we compute the Mahalanobis distance between them as:
-$D_{M}(\mathbf{x})=\sqrt{(\mathbf{x}-\boldsymbol{\mu})^{T}\Sigma^{-1}(\mathbf{x}-\boldsymbol{\mu})}$ (25)
-where $\mathbf{x}$ is the vector of observed societal metrics for a given type, $\boldsymbol{\mu}$ is the mean vector of another societal type, and $\Sigma$ is the covariance matrix of the distribution of human behavior.
-To complement the Mahalanobis distance, we compute the Chi-Squared p-value, which indicates the statistical significance of the observed differences.
-Since the squared Mahalanobis distance follows a Chi-Squared distribution with degrees of freedom equal to the number of dimensions, the p-value provides a probabilistic interpretation of the similarity between societal types.
-Lower p-values suggest significant differences between the compared distributions, while higher p-values indicate that the societal structures are statistically indistinguishable.
-This methodology enables us to identify when two societies exhibit fundamentally different organization patterns and to characterize the extent of these differences.
-Appendix F Experiment A
-Additional implementation details and results from the first experiment reported in the main paper:
-  • We use the same JHG parameter settings as was used by Skaggs et al.
-[41].
-That is, $\alpha=0.2$, $\beta=0.5$, $c_{\rm give}=1.3$, $c_{\rm keep}=0.95$, $c_{\rm take}=1.6$.
-Furthermore, each agent had $2*|I|$ tokens to allocate in each round.
-  • eCABs were evolved with a single set of genes (rather than the 3 sets used by Skaggs et al.
-[41]) and were trained for 200 generations with variable initial probabilities.
-Simulation results were conducted with randomly selected parameterizations from the 100 parameterizations that were in the final generation.
-  • Additional precision along with standard for average economic actions in each society are provided in Table 4
-  • No GPU was used in the study.
-No high-powered computational equipment were used (only standard computing equipment with multiple cores).
-Training using the PSO algorithm was done on a 32-core machine, which took about 3 hours per run.
-Training using the EPDM algorithm took about 30 minutes on a machine with 16 cores.
-Agent-Society % Give % Keep % Take
+Table 1. Metrics used to compare the population dynamics of human and agent societies.
+Category Description Metrics
+Wealth & Power Wealth and influence of society members Mean Popularity, Gini Index
+Economic Behavior High-level actions (give, take, keep) taken by individuals in society % Give, % Take, % Keep, Evolution Coefficient
+Mixing Patterns Connections across society, including dyadic, triadic, and group relationships Reciprocity, Density, Entropy, Polarization
+Summary Comparison to humans across all metrics Mahalanobis distance
+4. Evaluating population dynamics
+Our goal is to model human behavior in the JHG so as to satisfy two criteria.
+First, the population dynamics of agent collectives should mirror the population dynamics of human societies.
+Second, individual agents should, ideally, also exhibit human-like behavior.
+In this section, we evaluate four models, created by combining parameter-based models (hTFT and hCAB111hCAB and hTFT denote parameterizations of CAB and TFT models, respectively, learned from human data.) with modeling algorithms (PSO and EPDM), with respect to the first criteria.
+In Section 5, we address the second criteria.
+We note that human behavior is likely impacted by group size.
+Understanding human behavior given all group sizes is valuable.
+However, given that our data sets (Section 4.2) consist of small-scale human groups (5-12 individuals), we only consider groups of this size in the evaluations made in this paper.
+While we anticipate that human behavior in larger groups could be modeled with the same algorithmic mechanisms, we leave such evaluations to future work.
+4.1. Experiment Design
+We simulate societies of hTFT-PSO, hTFT-EPDM, hCAB-PSO, and hCAB-EPDM agents in the JHG, and compare their dynamics to those of games played by human players.
+We also consider two baseline agents: (1) Random, which randomly allocates tokens with the percentages of give, take, and keep that match the percentages observed in the training set (Section 4.2) and (2) eCAB agents (skaggs2024fostering).
+eCABs are CAB agents with parameterizations learned via evolutionary simulations rather than from human data.
+eCABs display somewhat effective behavior in the JHG, though their strategies are distinct from those used by humans skaggs2024fostering.
+To evaluate these agents, we compare their population dynamics with those of human players observed in the 15 games in our test set (Section 4.2).
+Simulations for agent populations were conducted with identical settings (initial popularities, number of agents, and game length) as games found in the test set.
+For each game in the test set, four simulations of each agent population were conducted with identical numbers of agents and rounds, bringing the total number of simulations for each agent to 60.
+In setting up experiments in this way, we can directly compare human and agent societies.
+For simulations with hTFT-EPDM, hCAB-EPDM, and eCAB agents, parameterizations were randomly selected from the learned pool of strategies for each game.
+On the other hand, for hTFT-PSO and hCAB-PSO, the parameterizations to which the PSO algorithm converged in four separate runs of the PSO algorithm were used.
+4.2. Data Sets
+Our data set, which is provided in the supplementary material, consists of 66 games played primarily by experienced human players.
+47 of these 66 games had only human players.
+The other 19 games were played by both human and bot players.
+Of the 47 all-human games, we randomly selected 15 of these games for the test set.
+The remaining 51 games were designated as training data.
+Games in the training set had 5-12 players with an average of 8.08 players (of which 6.57 were humans).
+Across these games, we have a total of 335 human player games (which comprise $\Gamma$).
+The average number of rounds in these player games was 24.96.
+Games in the test set were played by 6-11 human players (average was 7.93 players).
+Games with more than 30 rounds were truncated at 30 rounds.
+4.3. Metrics to Evaluate Population Dynamics
+We evaluate population dynamics using the metrics summarized in Table 1.
+These metrics are grouped into four categories, each designed to evaluate a particular aspect of population dynamics.
+Metrics in the first category quantify the wealth and power wielded by society members.
+We use Mean Popularity, an approximation of Katz centrality (katz1953new), and Gini Index, which measures the distribution of wealth across society.
+Economic behavior summarizes player actions across society, including the percentages of token allocations of each type (keep, take, and give) and the average change in player token allocations from round to round (Evolution Coefficient).
+Next, we assess societal mixing patterns, or the way that players both join together in pairs and groups, and fragment and segregate.
+These mixing patterns are quantified using two measures of reciprocity, density, entropy, and polarization.
+Finally, we summarize the dissimilarity between agent and human populations across all these metrics using Mahalanobis distance (mahalanobis1936generalized).
+Formal definitions of each metric, as well as motivation for why we chose these (as opposed to other) metrics, are provided in SM-5.
+4.4. Results
+Figure 2 displays measures of human and agent population dynamics with respect to individual metrics. We discuss each by category.
+(a)                  (b)                  (c)
+Society % Give % Keep % Take
 Humans $74.9\pm 1.0$ $18.8\pm 0.7$ $6.2\pm 0.8$
 Random $72.0\pm 0.0$ $24.0\pm 0.0$ $4.0\pm 0.0$
-hTFT-PSO $97.3\pm 1.8$ $2.7\pm 1.8$ $0.0\pm 0$ 0.0
+hTFT-PSO $97.3\pm 1.8$ $2.7\pm 1.8$ $0.0\pm 0.0$
 hTFT-EPDM $69.9\pm 0.3$ $14.9\pm 0.8$ $15.2\pm 0.4$
-eCAB $94.7\pm 1.8$ $5.2\pm 0.0$ $0.0\pm 0$ 0.0
-hCAB-PSO $100.0\pm 0.0$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
+eCAB $94.7\pm 1.8$ $5.2\pm 0.0$ $0.0\pm 0.0$
+hCAB-PSO $100.0\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$
 hCAB-EPDM $73.6\pm 1.0$ $19.1\pm 0.4$ $7.3\pm 0.6$
-Table 4: Mean percent of each economic actions in human and agent societies. $\pm$ values show show standard error of the mean.
-Appendix G Experiment B: User study details
-Additional information about the second experiment (user study) reported in the main paper:
-  • The user study was approved by the authors’ institutional review board (it was given exempt status). Participants gave consent to participated in the study through the approved process.
-  • Participants were paid approximately \$16 (USD) per hour while participating in the study.
-A high-score list (based on final popularity in each game) was kept to motivate participants to perform well.
-  • Games were played using the publicly available online platform published by Skaggs et al.
-[41] (permission was obtained from the authors of that work to use the platform for this study).
-This online platform can be found at juniorhighgame.com.
-  • Story plots for all eight games played in the user study are provided below.
+[…]
+Figure 2.
+Comparisons of human and agent societies.
+Dots (a-c), numbers (d), and lines (e-f) denote the mean, error ovals (which do not account for covariance), $\pm$, and error ribbons show the standard error.
+Wealth and Power: Figure 2a compares agent and human populations with respect to Mean Popularity and Gini Index.
+In these two regards, hCAB-EPDM populations most closely mirror those of humans.
+Random populations tend to have both low Gini Index and low mean popularity.
+hTFT-EPDM populations tend to have slightly higher Gini Index and lower popularity, while hTFT-PSO, hCAB-PSO, and eCAB populations all have higher Mean Popularity and lower Gini Index.
+While high Mean Popularity and lower Gini Index is desirable (though this does not necessarily indicate robust behavior in general), our objective is to identify models that mimic human behavior.
+With respect to Mean Popularity and Gini Index, the CAB model coupled with EPDM parameter modeling results in the model that most closely reflects human population dynamics.
+Economic Behavior: In the JHG, wealth outcomes are the result of economic behavior.
+eCAB agents and algorithms that model parameters using PSO all have higher Mean Popularity and lower Gini Index than human populations.
+These agents allocate most of their tokens to giving.
+They keep and take little, if any (Figure 2d).
+On the other hand, humans and agents modeled with EPDM have higher variability in strategies across players and tend to keep and take more tokens.
+This produces greater societal inequality and less overall prosperity.
+As such, agents produced using EPDM have token allocation distributions that closer to those of humans.
+The Evolution Coefficient measures the consistency of player token allocations over rounds.
+Figure 2e shows that none of the agent populations display similar behavior to humans in this regard.
+CAB agents are too consistent in their allocations (though hCABs are less consistent than eCABs), showing less change in token allocations from round to round than humans.
+On the other hand, TFT agents display a sort of multiple personality disorder.
+The act of reciprocating actions from the previous round produces oscillating changes in token allocations that are clearly not human-like.
+However, the differences in token allocations of TFT agents over two rounds are quite small.
+Mixing Patterns: We measure two forms of reciprocity: Immediate Reciprocity (reciprocation from one round to the next) and Overall Reciprocity (reciprocation over a complete game).
+In the JHG, humans have moderate levels of both forms of reciprocity (Figure 2b).
+EPDM agents exhibit similar levels of reciprocity, with hCAB-EPDM’s rates being only slightly higher than people’s.
+On the other hand, eCABs and PSO reciprocation rates are far higher than those of the humans in evaluated in our test set.
+The density and entropy measures, shown in Figure 2c, provide information about the number (as measured by Density) and strength (as measured by Entropy–this metric measures whether all trading partners are valued similarly) of trading partners.
+In these regards, hCAB-EPDM agents once again tend to be very similar to those of human players.
+hCAB-PSO agents have similar Density, but have higher Entropy.
+On the other hand, agents modeled using the TFT model tend to be less similar to humans in both regards.
+Thus, hCAB agents mirror these human behaviors better than TFT agents, with hCAB-EPDM being the closest.
+Finally, we consider Polarization (Figure 2f), which describes the degree to which players segregate into groups.
+Since the CAB algorithm is based on group formation, while TFT is not, it is not surprising to observe that CAB agents display higher Polarization than TFT agents.
+These higher levels are similar to the Polarization measures we observe in human games in the test set.
+This again shows that the CAB model can be tuned to better model human dynamics than the TFT (matching) model.
+Table 2.
+Mahalanobis distance (lower is better) between human and agent populations over all metrics.
+$p>0.05$ indicates that we cannot reject the null hypothesis (that they come from the same distribution).
+Algorithm Mahalanobis Distance Chi-Square Test
+Random 18.822 $p<0.001$
+eCAB 7.089 $p<0.001$
+hTFT-PSO 11.565 $p<0.001$
+hTFT-EPDM 7.656 $p<0.001$
+hCAB-POS 9.216 $p<0.001$
+hCAB-EPDM 3.393 $\mathbf{p=0.486}$
+Summary Metric: Table 2 gives the Mahalanobis distance between the various agent populations and our human population across the eleven individual metrics.
+Of the six agent populations, hCAB-EPDM populations most closely mirror human societies from our test set.
+Indeed, a Chi-Square test comparing human and hCAB-EPDM populations produces a p-value ($p=0.486$) indicating that we cannot reject the null hypothesis (that they come from the same distribution), thus indicating a somewhat close fit across these measures.
+All other agent populations were statistically distinct from human populations ($p<0.001$ in all cases).
+These results show that the CAB algorithm paired with EPDM parameterization modeling produces agent populations whose dynamics are surprisingly close (with some variation–notably the Evolution Coefficient) to those of human populations in the JHG (first criteria).
+Is this relatively similar behavior limited to population dynamics, or does it also carry over into individual behaviors?
+In the next section, we complement the quantitative comparisons of human and agent populations with a user study designed to evaluate whether individual hCAB-EPDM agents plausibly exhibit human behavior (second criteria).
+5. Interacting with human players
+To further illustrate the behavior of hCAB-EPDM agents (hereafter referred to as hCABs), we conducted a user study in which experienced human players interacted with hCABs in the JHG.
+From this study, we (1) illustrate the behavior of individual hCABs when interacting with humans and (2) quantify how well human participants were able to distinguish hCABs from humans.
+5.1. User Study Design
+We recruited eight people, each of which had prior experience playing the JHG, to participate in the study.
+The study proceeded in a series of four periods.
+In each period, the participants were randomly divided in half.
+Each group of four participants was paired with four (randomly selected) hCABs to form two games played by eight players each.
+Thus, eight games were played across the four periods.
+Games lasted for 21, 20, 17, and 15 rounds in Periods 1-4, respectively (players were not told game lengths).
+For consistency, only the first 15 rounds of each game were used in our data analysis.
+Participants were asked to play so as to become as popular as possible by the end of each game, while also identifying whether associates were humans or bots.
+Participants were told in advance that each game was composed of four people and four bots.
+After each game, participants completed a survey asking them to identify each associate as a human or a bot.
+We also compare humans and hCABs with respect to their popularity and the percentage of tokens used to give, keep, and take.
+5.2. Results
+Figure 3. Mean popularity over time of humans and hCABs in the user study. Error ribbons show the standard error.
+[…]
+A comparison of the mean popularity over time of humans and bots in the user study.
+Table 3. The percentage of tokens allocated to giving, keeping, and taking by humans and hCABs in the user study.
+Society % Give % Keep % Take
+Humans 60 25 15
+hCABs 75 20 5
+Table 4. The percentage of players correctly identified as human or bot by human participants verses random chance.
+Guesser % Correct Designation
+Humans 56.7
+Chance 57.1
+Figure 4.
+Illustrative game showing the popularity of players over time (top) and network structure in select rounds (bottom).
 In the graphs, arrows indicate token transactions in the current round (green = give; red = take), while nodes more connected historically tend to be closer to each other.
+[…]
+Illustration of a human-bot game showing popularity of the players over time as well as relationship graphs in select rounds.
+Over the eight games, hCABs had slightly higher average popularity than humans in rounds 5-15 (Figure 3).
+This was potentially caused by some human participants being quite aggressive, attacking substantially more than hCABs (Table 3).
+As attacks sometimes lead to retaliation and prolonged fighting, these players often subsequently lost popularity.
+While the distribution of human actions was different than humans in our data sets (see previous section), average hCAB behavior remained about the same.
+Despite these differences between these particular human participants and hCABs, participants were unable to distinguish each other from hCABs on average.
+Table 4 shows that participants correctly identified whether their associates were human or bot 56.7% of the time.
+Given that participants knew that four of their associates were bots and three were humans, random chance would have produced a correct detection rate of 57.1%.
+Therefore, these participants were, on average, no better than random in identifying whether their associates were humans or hCABs.
+Viewing hCAB behavior in individual games provides further evidence of this claim.
+For example, consider Figure 4, which depicts one of the eight games from the study (similar graphs for the other games are provided in SM-7).
+In Round 9, two human players (p1 and p2) attacked each other.
+Interestingly, three of the four hCABs and the two other humans all joined with p7 in attacking p1 in Round 10 (which greatly weakened p1 for the remainder of the game).
+That the hCABs all chose to take the same side in the dispute between p7 and p1 as the other human players (siding with the better connected individual) illustrates the ability of hCABs to reason about groups and power in a way that led them to mirror human behavior in this situation.
+Similar observations can be made about how hCABs joined together in subgroups with humans and other hCABs in these games.
+These results give credence to the argument that the behaviors of individual hCABs do seem human-like.
+Similarities between humans and hCABs can be seen in both population dynamics and in individual behaviors.
+However, we caution against over-extrapolating these results.
+The user study measured only a small set of experienced (eight) human players.
+Differences could emerge in a larger user study.
+Furthermore, population dynamics (see the previous section) do show differences between humans and hCABs.
+That said, these results do, by and large, indicate that hCABs are surprisingly difficult to distinguish from humans.
+6. Conclusions and delimitations
+In this paper, we compared and contrasted several algorithms for modeling human behavior in a strategic network game called the Junior High Game (JHG).
+We considered approaches that differ with respect to the underlying assumptions of human behavior they use (behavior matching verses community-aware behavior) and the statistical moments they model (mean verses distribution).
+Results indicate that a parameter-based community-aware model (CAB (skaggs2024fostering)) coupled with evolutionary population distribution modeling (EPDM) can produce agents that closely reflect human population dynamics in the JHG (albeit with differences).
+Furthermore, results of a user study indicate that the individual behaviors of these agents are difficult to distinguish from those of humans.
+The CAB model consists of a variety of hand-coded reasoning functions modulated by thirty parameters.
+Such a modeling approach has advantages and disadvantages compared to a more tabula rasa approach (such as a graph neural network).
+One advantage is that an effective model of human behavior can be derived using only a small data set.
+This is valuable, since obtaining large amounts of quality data in strategic network games can be challenging.
+A second advantage is that the trained strategies are interpretable, thus potentially shedding insight on human behavior.
+On the other hand, a tabula rasa approach could potentially learn other kinds of strategies, and is therefore a potentially promising area for future work once larger data sets are obtained.
+hCAB-EPDM provides a baseline for evaluating such future methods.
+While our results indicate that hCAB-EPDM agents are, perhaps surprisingly, effective models of human behavior in the JHG, we caution against overly broad extrapolations.
+For example, the groups modeled in this paper were small (6-11 players).
+Larger societies have different characteristics.
+Additionally, the hCAB-EPDM agents used in this paper were trained on a small data set.
+Given the complex state and strategy space of the JHG, other human players (from, for example, different cultures) could potentially use different strategies not found in this data set.
+Though the CAB algorithm can be configured (via its parameters) to model many different kinds of behavior, it is possible that other communities of human players could display vastly different norms of acceptable behavior that may not be effectively modeled by parameterizations of the CAB algorithm.
+Finally, the value of the results described in this paper somewhat hinges on the degree to which the strategic network game we selected (the JHG) models human networks.
+While the JHG models important properties of such networks, the world is certainly more complex than any game can model.
+Future work should consider refinements and enhancements to study other complexities of human networks.
+These and other future works can potentially help us to continue to better understand and model human networks.
+7. Supplementary Material (SM)
+Supplementary documentation, results, and code are available at: https://github.com/jakecrandall/AAMAS2026-hCABs.git